### v1

# Noncommutativepre-Poisson bialgebras and relative Rota-Baxter operators

Abstract.
In this paper, we develop the bialgebra theory for coherent noncommutative pre-Poisson algebras and establish equivalences among matched pairs, Manin triples, the phase space of noncommutative Poisson algebras and noncommutative pre-Poisson bialgebras. The investigation of coboundary noncommutative pre-Poisson bialgebras naturally leads to the noncommutative pre-Poisson Yang-Baxter equation (NPP-YBE). We prove that a symmetric solution of the NPP-YBE gives rise to a (coboundary) noncommutative pre-Poisson bialgebra. Moreover, we demonstrate how solutions without the symmetry condition can also generate such bialgebras. This motivates the introduction of quasi-triangular and factorizable noncommutative pre-Poisson bialgebras. In particular, we show that a solution of the NPP-YBE with an invariant skew-symmetric part yields a quasi-triangular noncommutative pre-Poisson bialgebra. Such solutions are further interpreted as relative Rota-Baxter operators with weights. Finally, we establish a one-to-one correspondence between quadratic Rota-Baxter noncommutative pre-Poisson algebras and factorizable noncommutative pre-Poisson bialgebras.
Key words and phrases:
noncommutative pre-Poisson algebra, noncommutative pre-Poisson Yang-Baxter equation, quasi-triangular noncommutative pre-Poisson bialgebra, factorizable noncommutative pre-Poisson bialgebra, relative Rota-Baxter operator
2020 Mathematics Subject Classification:
17A30, 17B63, 17B38, 16T10,17B62
1. Introduction
The purpose of this paper is to develop a bialgebra theory for noncommutative pre-Poisson algebras. The study of coboundary noncommutative pre-Poisson bialgebras naturally gives rise to the noncommutative pre-Poisson Yang-Baxter equation (NPP-YBE). In particular, we investigate quasi-triangular and factorizable noncommutative pre-Poisson bialgebras, which constitute a subclass of coboundary noncommutative pre-Poisson bialgebras. Furthermore, we explicitly construct a correspondence linking factorizable noncommutative pre-Poisson bialgebras to quadratic Rota-Baxter noncommutative pre-Poisson algebras.
1.1. (Noncommutative) Poisson algebras and (noncommutative) pre-Poisson algebras
A Poisson algebra is an algebraic structure that combines a Lie algebra with a commutative associative algebra, linked together by the Leibniz rule. Poisson algebras originated from the study of Hamiltonian mechanics and have since become important in many areas of mathematics and mathematical physics. These include Poisson geometry [28, 31], algebraic geometry [11, 23], quantum mechanics [8], quantum groups [9] and quantization theory [13, 14].
Aguiar integrated the structures of Zinbiel algebras [20] and pre-Lie algebras [5] on the same vector space, thereby introducing the notion of a pre-Poisson algebra [1]. Zinbiel algebras are Koszul dual to Leibniz algebras, and they are special dendriform algebras. The anti-commutator of a Zinbiel algebra gives a commutative associative algebra, and the commutator of a pre-Lie algebra gives a Lie algebra. A pre-Poisson algebra induces a Poisson algebra naturally through the sub-adjacent commutative associative algebra of the Zinbiel algebra and the sub-adjacent Lie algebra of the pre-Lie algebra. Conversely, a Rota-Baxter operator (more generally an $\mathcal{O}$-operator action) on a Poisson algebra gives rise to a pre-Poisson algebra. We can summarize these relations by the following diagram:
Zinbiel algebra + pre-Lie algebra sub-adjacent  Pre-Poisson algebrasub-adjacent  comm associative algebra + Lie algebra   Rota-Baxter  operator Poisson algebra.   Rota-Baxter  operator
A noncommutative Poisson algebra, first formally introduced by Xu [32], provides a natural framework for various geometric contexts. It extends the classical notion of a Poisson algebra by replacing the underlying commutative associative algebra with a general (not necessarily commutative) associative algebra, while retaining a compatible Lie algebra structure linked via the Leibniz rule. This relaxation of commutativity results in significantly richer and more complex algebraic behaviors. Noncommutative Poisson algebras have been studied by many authors from different aspects, see [15, 16, 17, 33, 19] and reference therein.
Definition 1.1.
A noncommutative Poisson algebra is a triple $(A,\cdot,[\ ,\ ])$ such that $(A,\cdot)$ is an associative algebra (not necessarily commutative), (A,[,])(A,[\ ,\ ]) is a Lie algebra and the following Leibniz rule holds:
(1) $x,y\cdot z]=[x,y]\cdot z+y\cdot[x,z],\quad\forall~x,y,z\in A.$
It is called coherent if, in addition,
(2) $x,y\cdot z]+[y,z\cdot x]+[z,x\cdot y]=0,\quad\forall~x,y,z\in A.$
Since Zinbiel algebras can be viewed as commutative dendriform algebras [21], a natural generalization is to combine dendriform algebras with pre-Lie algebras. This leads to the notion of a noncommutative pre-Poisson algebra [19]. The correspondence between Poisson algebras and pre-Poisson algebras is formally analogous to that between their noncommutative versions, a relationship illustrated in the following diagram,
dendriform algebra + pre-Lie algebra sub-adjacent  noncomm pre-Poisson algebrasub-adjacent  associative algebra + Lie algebra   Rota-Baxter  operatornoncomm Poisson algebra  Rota-Baxter  operator
1.2. Quasi-triangular bialgebras
A bialgebraic structure comprises an algebra and a coalgebra equipped with certain compatibility conditions. In the early 1980s, Drinfeld established the theory of Lie bialgebras [10], which were found to have deep connections with the classical Yang-Baxter equation and classical integrable systems. Subsequently, V. Zhelyabin introduced associative D-bialgebras in [34, 35]. Aguiar later in [2] studied antisymmetric infinitesimal bialgebras as the associative analogue of Lie bialgebras, which are also equivalent to double constructions of Frobenius algebras [4]. Following the development of infinitesimal bialgebra theory [2, 4], similar bialgebraic frameworks have been extended to various other algebraic structures, such as pre-Lie algebras [3], Leibniz algebras [27] and Novikov algebras [12]. A Manin triple of Poisson algebras corresponds to a Poisson bialgebra [22], offering a natural framework for constructing compatible Poisson brackets in integrable systems. Unlike in the commutative case, a direct generalization of Poisson bialgebra theory is not straightforward due to fundamental differences in representation theory. In [19], the authors developed the noncommutative Poisson bialgebra theory using quasi-representations and the corresponding cohomology theory [7, 33].
Within the theory of Lie bialgebras, coboundary Lie bialgebras especially quasi-triangular Lie bialgebras, which play a fundamental role in mathematical physics. As a specialized subclass of quasi-triangular Lie bialgebras, factorizable Lie bialgebras [18] provide a crucial connection between classical $r$-matrices and certain factorization problems. They exhibit diverse applications in integrable systems, see [24, 25] and references therein. Recently, these results on factorizable and quasi-triangular structures have been successfully extended to antisymmetric infinitesimal bialgebras [26], pre-Lie bialgebras [29] and Leibniz bialgebras [6].
1.3. Outline of the paper
A bialgebra theory has been developed for the usual (commutative) pre-Poisson algebras, so-called pre-Poisson bialgebras [36], in terms of the representation theory of pre-Poisson algebras. However, a direct generalization of this framework does not apply to noncommutative pre-Poisson algebras. This limitation arises because the dual of the regular representation of a noncommutative Poisson algebra is generally not a representation, but only a quasi-representation. To overcome this obstacle, we develop a bialgebra theory specifically for coherent noncommutative pre-Poisson algebras. Since every pre-Poisson algebra is in particular a coherent noncommutative pre-Poisson algebra, such a generalization is both natural and well justified.
This paper is organized as follows. In Section 2, we introduce quasi-representations and matched pairs for noncommutative pre-Poisson algebras. The notion of a phase space of a noncommutative Poisson algebra is also studied, which later be used to characterize the bialgebra theory. Section 3 is devoted to establishing a bialgebra theory for noncommutative pre-Poisson algebras. Investigating the coboundary case leads to the noncommutative pre-Poisson Yang-Baxter equation (NPP-YBE), whose symmetric solutions give rise to noncommutative pre-Poisson bialgebras. We further define $\mathcal{O}$-operators on noncommutative pre-Poisson algebras and use them to construct symmetric solutions of the NPP-YBE. In Section 4, we explore quasi-triangular and factorizable noncommutative pre-Poisson bialgebras. We prove that the double of any noncommutative pre-Poisson bialgebra naturally carries a factorizable structure. Finally, Section 5 presents the notion of quadratic Rota-Baxter noncommutative pre-Poisson algebras and develops a one-to-one correspondence between these and factorizable noncommutative pre-Poisson bialgebras.
[…]
$r_{12}\ast r_{13}:=\sum_{i,j}a_{i}\ast a_{j}\otimes b_{i}\otimes b_{j},\;r_{23}\ast r_{12}:=\sum_{i,j}a_{j}\otimes a_{i}\ast b_{j}\otimes b_{i},\;r_{31}\ast r_{23}:=\sum_{i,j}b_{i}\otimes a_{j}\otimes a_{i}\ast b_{j},$
$r_{21}\ast r_{13}:=\sum_{i,j}b_{i}\ast a_{j}\otimes a_{i}\otimes b_{j},\;r_{32}\ast r_{21}:=\sum_{i,j}b_{j}\otimes b_{i}\ast a_{j}\otimes a_{i},\;r_{31}\ast r_{32}:=\sum_{i,j}b_{i}\otimes b_{j}\otimes a_{i}\ast a_{j},$
$r_{13}\ast r_{32}:=\sum_{i,j}a_{i}\otimes b_{j}\otimes b_{i}\ast a_{j},\;r_{23}\ast r_{21}:=\sum_{i,j}b_{j}\otimes a_{i}\ast a_{j}\otimes b_{i},\;r_{21}\ast r_{31}:=\sum_{i,j}b_{i}\ast b_{j}\otimes a_{i}\otimes a_{j},$
$r_{23}\ast r_{13}:=\sum_{i,j}a_{i}\otimes a_{j}\otimes b_{i}\ast b_{j},\;r_{12}\ast r_{31}:=\sum_{i,j}a_{i}\ast b_{j}\otimes b_{i}\otimes a_{j}.$
2. (Quasi-) representations and phase spaces of noncommutative Poisson algebras
In this section, we introduce the notions of (quasi-) representations, matched pairs and Manin triples for noncommutative pre-Poisson algebras. We also consider the phase space of noncommutative Poisson algebras. Finally, we exhibit the relationship between matched pairs, Manin triples of noncommutative pre-Poisson algebras and the phase space of noncommutative Poisson algebras. We begin by recalling the definitions of pre-Lie algebras [5], dendriform algebras [21] and noncommutative pre-Poisson algebras [19].
2.1. (Quasi-) representations of noncommutative pre-Poisson algebras
Let $A$ be a vector space with a bilinear product $\ast:A\otimes A\longrightarrow A$, $(A,\ast)$ is called a pre-Lie algebra if
$(x\ast y)\ast z-x\ast(y\ast z)=(y\ast x)\ast z-y\ast(x\ast z),~\forall~x,y,z\in A.$
Definition 2.1.
A dendriform algebra is a vector space $A$ together with two bilinear maps $\succ,\prec:A\otimes A\longrightarrow A$ such that the following equalities hold for all $x,y,z\in A$:
$(x\prec y)\prec z=x\prec(y\succ z+y\prec z),$
$(x\succ y)\prec z=x\succ(y\prec z),$
$x\succ(y\succ z)=(x\succ y+x\prec y)\succ z.$
Definition 2.2.
A noncommutative pre-Poisson algebra is a quadruple $(A,\succ,\prec,\ast)$ such that $(A,\succ,\prec)$ is a dendriform algebra and $(A,\ast)$ is a pre-Lie algebra satisfying the following compatibility conditions:
(3) $(x\ast y-y\ast x)\succ z$ $=$ $x\ast(y\succ z)-y\succ(x\ast z),$
(4) $x\prec(y\ast z-z\ast y)$ $=$ $y\ast(x\prec z)-(y\ast x)\prec z,$
(5) $(x\succ y+x\prec y)\ast z$ $=$ $(x\ast z)\prec y+x\succ(y\ast z).$
A noncommutative pre-Poisson algebra $(A,\succ,\prec,\ast)$ is called coherent if it also satisfies
(6) $(x\succ y+x\prec y)\ast z=x\ast(y\succ z)+y\ast(z\prec x).$
Let $(A,\succ,\prec,\ast)$ be a (coherent) noncommutative pre-Poisson algebra. Define
$x\cdot y=x\succ y+x\prec y,\ \ \ [x,y]=x\ast y-y\ast x,~\forall~x,y\in A.$
Then $(A,\cdot,[\ ,\ ])$ is a (coherent) noncommutative Poisson algebra, which is called the sub-adjacent noncommutative Poisson algebra of $(A,\succ,\prec,\ast)$.
Example 2.3.
Let $A$ be a 2-dimensional vector space with a basis $\{e_{1},e_{2}\}$. Define bilinear maps $\succ,\prec,\ast:A\otimes A\longrightarrow A$ respectively by (only non-zero multiplications are listed):
$e_{1}\succ e_{2}=e_{2},\ \ \ e_{1}\succ e_{1}=e_{1},\ \ \ e_{1}\prec e_{2}=-e_{2},$
$e_{2}\prec e_{1}=e_{2},\ \ \ e_{1}\ast e_{1}=e_{1},\ \ \ e_{2}\ast e_{1}=e_{2}.$
By a direct computation, $(A,\succ,\prec,\ast)$ is a coherent noncommutative pre-Poisson algebra.
Example 2.4.
Let $A$ be a 3-dimensional vector space with a basis $\{e_{1},e_{2},e_{3}\}$. Define binary products $\succ,\prec,\ast:A\otimes A\longrightarrow A$ respectively by (only non-zero multiplications are listed):
$e_{2}\succ e_{2}=e_{3},\ \ \ e_{2}\prec e_{2}=-e_{3},\ \ \ e_{1}\ast e_{2}=e_{3},\ \ \ e_{2}\ast e_{2}=e_{1}.$
By a direct computation, $(A,\succ,\prec,\ast)$ is a coherent noncommutative pre-Poisson algebra.
We now turn to the study of quasi-representations and representations of noncommutative pre-Poisson algebras.
Definition 2.5.
Let $(A,\succ,\prec,\ast)$ be a noncommutative pre-Poisson algebra, $V$ a vector space and $l_{\succ},r_{\succ},l_{\prec},r_{\prec},l_{\ast},r_{\ast}:A\longrightarrow\hbox{End}(V)$ be linear maps.
  1. (a)
    $(V,l_{\succ},r_{\succ},l_{\prec},r_{\prec},l_{\ast},r_{\ast})$ is called a quasi-representation of $(A,\succ,\prec,\ast)$ if $(V,l_{\succ},r_{\succ},l_{\prec},r_{\prec})$ is a representation of the dendriform algebra $(A,\succ,\prec)$, $(V,l_{\ast},r_{\ast})$ is a representation of the pre-Lie algebra $(A,\ast)$ and they satisfy the following conditions for all $x,y\in A$:
    (7) $l_{\succ}(x\ast y-y\ast x)=l_{\ast}(x)l_{\succ}(y)-l_{\succ}(y)l_{\ast}(x),$
    (8) $r_{\succ}(y)(l_{\ast}(x)-r_{\ast}(x))=l_{\ast}(x)r_{\succ}(y)-r_{\succ}(x\ast y),$
    (9) $l_{\prec}(x)(l_{\ast}(y)-r_{\ast}(y))=l_{\ast}(y)l_{\prec}(x)-l_{\prec}(y\ast x),$
    (10) $r_{\prec}(y\ast x-x\ast y)=l_{\ast}(y)r_{\prec}(x)-r_{\prec}(x)l_{\ast}(y),$
    (11) $r_{\ast}(y)(l_{\succ}(x)+l_{\prec}(x))=l_{\prec}(x\ast y)+l_{\succ}(x)r_{\ast}(y),$
    (12) $r_{\ast}(x)(r_{\succ}(y)+r_{\prec}(y))=r_{\prec}(y)r_{\ast}(x)+r_{\succ}(y\ast x).$
  2. (b)
    A quasi-representation $(V,l_{\succ},r_{\succ},l_{\prec},r_{\prec},l_{\ast},r_{\ast})$ is called a representation of $(A,\succ,\prec,\ast)$ if the following conditions are satisfied:
    (13) $r_{\succ}(x)(l_{\ast}(y)-r_{\ast}(y))=l_{\succ}(y)r_{\ast}(x)-r_{\ast}(y\succ x),$
    (14) $l_{\prec}(x)(l_{\ast}(y)-r_{\ast}(y))=r_{\prec}(y)r_{\ast}(x)-r_{\ast}(x\prec y),$
    (15) $l_{\ast}(x\succ y+x\prec y)=r_{\prec}(y)l_{\ast}(x)+l_{\succ}(x)l_{\ast}(y).$
By Eqs. (13)-(15), we have
(16) $r_{\cdot}(x)(l_{\ast}(y)-r_{\ast}(y))+l_{\cdot}(y)(l_{\ast}(x)-r_{\ast}(x))=l_{\ast}(y\cdot x)-r_{\ast}(y\cdot x),$
(17) $r_{\cdot}(x)(l_{\ast}(y)-r_{\ast}(y))+(l_{\ast}(y)-r_{\ast}(y))r_{\cdot}(x)=r_{\cdot}(y\ast x-x\ast y).$
Let $A$ and $V$ be vector spaces. For a linear map $f:A\longrightarrow\hbox{End}(V)$, define a linear map $f^{*}:A\longrightarrow\hbox{End}(V^{*})$ by $\langle f^{*}(x)u^{*},v\rangle=\langle u^{*},f(x)v\rangle$ for all $x\in A,u^{*}\in V^{*},v\in V$, where $\langle\ ,\ \rangle$ is the usual pairing between $V$ and $V^{*}$.
Proposition 2.6.
Let $(A,\succ,\prec,\ast)$ be a noncommutative pre-Poisson algebra.
  1. (a)
    If $(V,l_{\succ},r_{\succ},l_{\prec},r_{\prec},l_{\ast},r_{\ast})$ is a quasi-representation of $(A,\succ,\prec,\ast)$, then $(V^{*},r_{\cdot}^{*},-l_{\prec}^{*},-r_{\succ}^{*},$ $l_{\cdot}^{*},r_{\ast}^{*}-l_{\ast}^{*},r_{\ast}^{*})$ is also a quasi-representation of $(A,\succ,\prec,\ast)$.
  2. (b)
    If $(V,l_{\succ},r_{\succ},l_{\prec},r_{\prec},l_{\ast},r_{\ast})$ is a quasi-representation and satisfies
    (18) $l_{\ast}(x)(x\succ y+x\prec y)=l_{\ast}(x)l_{\succ}(y)+l_{\ast}(y)r_{\prec}(x),$
    (19) $r_{\ast}(y)(l_{\succ}(x)+l_{\prec}(x))=l_{\ast}(x)r_{\succ}(y)+r_{\ast}(y\prec x),$
    (20) $r_{\ast}(x)(r_{\succ}(y)+r_{\prec}(y))=r_{\ast}(y\succ x)+l_{\ast}(y)l_{\prec}(x)),$
    then $(V^{*},r_{\cdot}^{*},-l_{\prec}^{*},-r_{\succ}^{*},l_{\cdot}^{*},r_{\ast}^{*}-l_{\ast}^{*},r_{\ast}^{*})$ is a representation of $(A,\succ,\prec,\ast)$.
  3. (c)
    If $(V,l_{\succ},r_{\succ},l_{\prec},r_{\prec},l_{\ast},r_{\ast})$ is a quasi-representation of $(A,\succ,\prec,\ast)$, then both $(V,l_{\succ},r_{\prec},-l_{\ast})$ and $(V,l_{\cdot},r_{\cdot},r_{\ast}-l_{\ast},r_{\ast})$ are quasi-representations of the sub-adjacent noncommutative Poisson algebra $(A,\cdot,[\ ,\ ])$, where $r_{\cdot}=r_{\succ}+r_{\prec},~l_{\cdot}=l_{\succ}+l_{\prec}$.
Proof.
(a) By Proposition 3.2.2 [4] and Proposition 3.3 [3], we know that $(V^{*},r_{\cdot}^{*},-l_{\prec}^{*},-r_{\succ}^{*},l_{\cdot}^{*})$ is a representation of $(A,\succ,\prec)$ and $(V^{*},r_{\ast}^{*}-l_{\ast}^{*},r_{\ast}^{*})$ is a representation of $(A,\ast)$. It therefore remains to verify that Eqs. (7)–(12) hold for the tuple $(V^{*},r_{\cdot}^{*},-l_{\prec}^{*},-r_{\succ}^{*},l_{\cdot}^{*},r_{\ast}^{*}-l_{\ast}^{*},r_{\ast}^{*})$. Indeed, Eq. (7) follows from Eqs. (7), (10) and (12); Eq. (8) follows from Eq. (9); Eq. (9) follows from Eq. (8); Eq. (10) follows from Eqs. (7) and (9); Eq. (12) follows from Eq. (11); Eq. (11) follows from Eq. (12).
(b) By Eq. (20), we get that Eq. (13) holds; by Eq. (19), we get that Eq. (14) holds; by Eqs. (19) and (20), we get that Eq. (15) holds.
(c) It can routinely be checked. ∎
Example 2.7.
Let $(A,\succ,\prec,\ast)$ be a noncommutative pre-Poisson algebra. Then
  1. (a)
    $(A,L_{\succ},R_{\succ},L_{\prec},R_{\prec},L_{\ast},R_{\ast})$ is a representation of $(A,\succ,\prec,\ast)$, which is called the regular representation of $(A,\succ,\prec,\ast)$.
  2. (b)
    $(A^{*},R_{\cdot}^{*},-L_{\prec}^{*},-R_{\succ}^{*},L_{\cdot}^{*},-\mathrm{ad}^{*},R^{*}_{\ast})$ is a quasi-representation of $(A,\succ,\prec,\ast)$.
  3. (c)
    $(A^{*},R_{\cdot}^{*},-L_{\prec}^{*},-R_{\succ}^{*},L_{\cdot}^{*},-\mathrm{ad}^{*},R^{*}_{\ast})$ is a representation of $(A,\succ,\prec,\ast)$ if and only if $(A,\succ,\prec,\ast)$ is coherent.
  4. (d)
    $(A,L_{\succ},R_{\prec},-L_{\ast})$ is a representation of the sub-adjacent noncommutative Poisson algebra $(A,\cdot,[\ ,\ ])$ and $(A^{*},R_{\prec}^{*},L_{\succ}^{*},-L_{\ast}^{*})$ is a quasi-representations of $(A,\cdot,[\ ,\ ])$.
  5. (e)
    $(A^{*},R_{\prec}^{*},L_{\succ}^{*},-L_{\ast}^{*})$ is a representation of the sub-adjacent noncommutative Poisson algebra $(A,\cdot,[\ ,\ ])$ if $(A,\succ,\prec,\ast)$ is coherent.
Let us recall the notion of matched pairs of noncommutative Poisson algebras, as introduced in [19].
Proposition 2.8.
Let $(P_{1},\cdot_{1},[\ ,\ ]_{1})$ and $(P_{2},\cdot_{2},[\ ,\ ]_{2})$ be two noncommutative Poisson algebras. Assume that there are linear maps $l_{1},r_{1},\rho_{1}:P_{1}\longrightarrow\hbox{End}(P_{2})$ and $l_{2},r_{2},\rho_{2}:P_{2}\longrightarrow\hbox{End}(P_{1})$. such that $(P_{1},P_{2},l_{1},r_{1},l_{2},r_{2})$ is a matched pair of associative algebras and $(P_{1},P_{2},\rho_{1},\rho_{2})$ is a matched pair of Lie algebras. Moreover, the following equalities hold for all $x,y\in P_{1},a,b\in P_{2}$,
(21) $\rho_{1}(x)(a\cdot_{2}b)=(\rho_{1}(x)a)\cdot_{2}b+a\cdot_{2}\rho_{1}(x)b-l_{1}(\rho_{2}(a)x)b-r_{1}(\rho_{2}(b)x)a,$
(22) $l_{1}(x)[a,b]_{2}=[a,l_{1}(x)b]_{2}-\rho_{1}(r_{2}(b)x)a-l_{1}(\rho_{2}(a)x)b+(\rho_{1}(x)a)\cdot_{2}b,$
(23) $\rho_{2}(a)(x\cdot_{1}y)=(\rho_{2}(a)x)\cdot_{1}y+x\cdot_{1}\rho_{2}(a)y-l_{2}(\rho_{1}(x)a)y-r_{2}(\rho_{1}(y)a)x,$
(24) $l_{2}(a)[x,y]_{1}=[x,l_{2}(a)y]_{1}-\rho_{2}(r_{1}(y)a)x-l_{2}(\rho_{1}(x)a)y+(\rho_{2}(a)x)\cdot_{1}y.$
Define two binary operations $\cdot$ and [,][\ ,\ ] on $P_{1}\oplus P_{2}$ by
$(x+a)\cdot(y+b)=x\cdot_{1}y+l_{\cdot_{2}}(a)y+r_{\cdot_{2}}(b)x+a\cdot_{2}b+l_{\cdot_{1}}(x)b+r_{\cdot_{1}}(y)a,$
$[x+a,y+b]=[x,y]_{1}+\rho_{2}(a)y-\rho_{2}(b)x+[a,b]_{2}+\rho_{1}(x)b-\rho_{1}(y)a.$
Then $(P_{1}\oplus P_{2},\cdot,[\ ,\ ])$ is a noncommutative Poisson algebra. Denote this noncommutative Poisson algebra by $A_{1}\bowtie A_{2}$ and $(P_{1},P_{2},l_{1},r_{1},\rho_{1},l_{2},r_{2},\rho_{2})$ satisfying the above conditions is called a matched pair of noncommutative Poisson algebras. Conversely, any noncommutative Poisson algebra that can be decomposed into a direct sum of two noncommutative Poisson subalgebras is obtained from a matched pair of noncommutative Poisson algebras.
Remark 2.9.
  1. (a)
    If the linear maps $l_{2},r_{2},\rho_{2}:P_{2}\rightarrow\text{End}(P_{1})$ are trivial, then the matched pair $(P_{1},P_{2},l_{1},r_{1},\rho_{1},l_{2},r_{2},\rho_{2})$ reduces to an A-noncommutative Poisson algebra. For the completeness, we list it in detail. Let $(P_{1},\cdot_{1},[\ ,\ ]_{1})$ and $(P_{2},\cdot_{2},[\ ,\ ]_{2})$ be two noncommutative Poisson algebras. Assume that there are linear maps $l_{1},r_{1},\rho_{1}:P_{1}\longrightarrow\hbox{End}(P_{2})$ such that $(P_{2},l_{1},r_{1},\rho_{1})$ is a representation of $(P_{1},\cdot_{1},[\ ,\ ]_{1})$ and the following compatible conditions hold for all $x,y\in P_{1},a,b\in P_{2}$:
    $l_{1}(x)(a\cdot_{2}b)=(l_{1}(x)a)\cdot_{2}b,\ \ \ \ \ r_{1}(x)(a\cdot_{2}b)=a\cdot_{2}(r_{1}(x)b),$
    $(r_{1}(x)a)\cdot_{2}b=a\cdot_{2}(l_{1}(x)b),\ \ \ \rho_{1}(x)[a,b]_{2}=[\rho_{1}(x)a,b]_{2}+[a,\rho_{1}(x)b]_{2},$
    $\rho_{1}(x)(a\cdot_{2}b)=(\rho_{1}(x)a)\cdot_{2}b+a\cdot_{2}\rho_{1}(x)b,\ \ \ l_{1}(x)[a,b]_{2}=[a,l_{1}(x)b]_{2}+(\rho_{1}(x)a)\cdot_{2}b.$
    Then $(P_{2},l_{1},r_{1},\rho_{1})$ is called an A-noncommutative Poisson algebra.
  2. (b)
    If the linear maps $l_{2},r_{2},\rho_{2}:P_{2}\rightarrow\text{End}(P_{1})$ and $(\cdot_{2},[\ ,\ ]_{2})$ are trivial, then the matched pair reduces to the semidirect product. Denote it simply by $A_{1}\ltimes A_{2}$.
Proposition 2.10.
Let $(A_{1},\succ_{1},\prec_{1},\ast_{1})$ and $(A_{2},\succ_{2},\prec_{2},\ast_{2})$ be two noncommutative pre-Poisson algebras. Suppose that there are linear maps $l_{\succ_{1}},r_{\succ_{1}},l_{\prec_{1}},r_{\prec_{1}},l_{\ast_{1}},r_{\ast_{1}}:A_{1}\longrightarrow\hbox{End}(A_{2})$ and $l_{\succ_{2}},r_{\succ_{2}},l_{\prec_{2}},r_{\prec_{2}},l_{\ast_{2}},r_{\ast_{2}}:A_{2}\longrightarrow\hbox{End}(A_{1})$. Define operations $\succ,\prec,\ast$ on $A_{1}\oplus A_{2}$ by
(25) $(x+a)\succ(y+b)=x{\succ_{1}}y+l_{\succ_{2}}(a)y+r_{\succ_{2}}(b)x+a{\succ_{2}}b+l_{\succ_{1}}(x)b+r_{\succ_{1}}(y)a,$
(26) $(x+a)\prec(y+b)=x{\prec_{1}}y+l_{\prec_{2}}(a)y+r_{\prec_{2}}(b)x+a{\prec_{2}}b+l_{\prec_{1}}(x)b+r_{\prec_{1}}(y)a,$
(27) $(x+a)\ast(y+b)=x{\ast_{1}}y+l_{\ast_{2}}(a)y+r_{\ast_{2}}(b)x+a{\ast_{2}}b+l_{\ast_{1}}(x)b+r_{\ast_{1}}(y)a,$
for all $x,y\in A_{1},a,b\in A_{2}$. Then $(A_{1}\oplus A_{2},\succ,\prec,\ast)$ is a noncommutative pre-Poisson algebra if and only if the following conditions hold:
  1. (a)
    $(A_{2},l_{\succ_{1}},r_{\succ_{1}},l_{\prec_{1}},r_{\prec_{1}},l_{\ast_{1}},r_{\ast_{1}})$ is a representation of $(A_{1},\succ_{1},\prec_{1},\ast_{1})$.
  2. (b)
    $(A_{1},l_{\succ_{2}},r_{\succ_{2}},l_{\prec_{2}},r_{\prec_{2}},l_{\ast_{2}},r_{\ast_{2}})$ is a representation of $(A_{2},\succ_{2},\prec_{2},\ast_{2})$.
  3. (c)
    $(A_{1},A_{2},l_{\succ_{1}},r_{\succ_{1}},l_{\prec_{1}},r_{\prec_{1}},l_{\succ_{2}},r_{\succ_{2}},l_{\prec_{2}},r_{\prec_{2}})$ is a matched pair of dendriform algebras.
  4. (d)
    $(A_{1},A_{2},l_{\ast_{1}},r_{\ast_{1}},l_{\ast_{2}},r_{\ast_{2}})$ is a matched pair of pre-Lie algebras.
  5. (e)
    The following compatible conditions hold for all $x,y\in A_{1},a,b\in A_{2}$:
    (28) $r_{\succ_{2}}(a)[x,y]_{1}=x\ast_{1}r_{\succ_{2}}(a)y+r_{\ast_{2}}(l_{\succ_{1}}(y)a)x-y\succ_{1}(r_{\ast_{2}}(a)x)-r_{\succ_{2}}(l_{\ast_{1}}(x)a)y,$
    (29) $(r_{\ast_{2}}(a)x-l_{\ast_{2}}(a)x)\succ_{1}y+l_{\succ_{2}}(l_{\ast_{1}}(x)a-r_{\ast_{1}}(x)a)y=x\ast_{1}(l_{\succ_{2}}(a)y)+r_{\ast_{2}}(r_{\succ_{1}}(y)a)x-l_{\succ_{2}}(a)(x\ast_{1}y),$
    (30) $(l_{\ast_{2}}(a)x-r_{\ast_{2}}(a)x)\succ_{1}y-l_{\succ_{2}}(r_{\ast_{1}}(x)a-l_{\ast_{1}}(x)a)y=l_{\ast_{2}}(a)(x\succ_{1}y)-x\succ_{1}(l_{\ast_{2}}(a)y)-r_{\succ_{2}}(r_{\ast_{1}}(y)a)x,$
    (31) $x\prec_{1}(r_{\ast_{2}}(a)y-l_{\ast_{2}}(a)y)+r_{\prec_{2}}(l_{\ast_{1}}(y)a-r_{\ast_{1}}(y)a)x=y\ast_{1}(r_{\prec_{2}}(a)x)+r_{\ast_{2}}(l_{\prec_{1}}(x)a)y-r_{\prec_{2}}(a)(y\ast_{1}x),$
    (32) $x\prec_{1}(l_{\ast_{2}}(a)y-r_{\ast_{2}}(a)y)+r_{\prec_{2}}(r_{\ast_{1}}(y)a-l_{\ast_{1}}(y)a)x=l_{\ast_{2}}(a)(x\prec_{1}y)-(l_{\ast_{2}}(a)x)\prec_{1}y-l_{\prec_{2}}(r_{\ast_{1}}(x)a)y,$
    (33) $l_{\prec_{2}}(a)(x\ast_{1}y-y\ast_{1}x)=x\ast_{1}(l_{\prec_{2}}(a)y)+r_{\ast_{2}}(r_{\prec_{1}}(y)a)x-(r_{\ast_{2}}(a)x)\prec_{1}y-l_{\prec_{2}}(l_{\ast_{1}}(x)a)y,$
    (34) $r_{\ast_{2}}(a)(x\cdot_{1}y)=(r_{\ast_{2}}(a)x)\succ_{1}y+l_{\prec_{2}}(l_{\ast_{1}}(x)a)y+x\succ_{1}(r_{\ast_{2}}(a)y)+r_{\succ_{2}}(l_{\ast_{1}}(y)a)x,$
    (35) $(r_{\cdot_{2}}(a)x)\ast_{1}y+l_{\ast_{2}}(l_{\cdot_{1}}(x)a)y=r_{\prec_{2}}(a)(x\ast_{1}y)+x\succ_{1}(l_{\ast_{2}}(a)y)+r_{\succ_{2}}(r_{\ast_{1}}(y)a)x,$
    (36) $(l_{\cdot_{2}}(a)x)\ast_{1}y+l_{\ast_{2}}(r_{\cdot_{1}}(x)a)y=(l_{\ast_{2}}(a)y)\prec_{1}x+l_{\prec_{2}}(r_{\ast_{1}}(y)a)x+l_{\succ_{2}}(a)(x\ast_{1}y),$
    (37) $r_{\succ_{1}}(x)[a,b]_{2}=a\ast_{2}r_{\succ_{1}}(x)b+r_{\ast_{1}}(l_{\succ_{2}}(b)x)a-b\succ_{2}(r_{\ast_{1}}(x)a)-r_{\succ_{1}}(l_{\ast_{2}}(a)x)b,$
    (38) $(r_{\ast_{1}}(x)a-l_{\ast_{1}}(x)a)\succ_{2}b+l_{\succ_{1}}(l_{\ast_{2}}(a)x-r_{\ast_{2}}(a)x)b=a\ast_{2}(l_{\succ_{1}}(x)b)+r_{\ast_{1}}(r_{\succ_{2}}(b)x)a-l_{\succ_{1}}(x)(a\ast_{2}b),$
    (39) $(l_{\ast_{1}}(x)a-r_{\ast_{1}}(x)a)\succ_{2}b-l_{\succ_{1}}(r_{\ast_{2}}(a)x-l_{\ast_{2}}(a)x)b=l_{\ast_{1}}(x)(a\succ_{2}b)-a\succ_{2}(l_{\ast_{1}}(x)b)-r_{\succ_{1}}(r_{\ast_{2}}(b)x)a,$
    (40) $a\prec_{2}(r_{\ast_{1}}(x)b-l_{\ast_{1}}(x)b)+r_{\prec_{1}}(l_{\ast_{2}}(b)x-r_{\ast_{2}}(b)x)a=b\ast_{2}(r_{\prec_{1}}(x)a)+r_{\ast_{1}}(l_{\prec_{2}}(a)x)b-r_{\prec_{1}}(x)(b\ast_{1}a),$
    (41) $a\prec_{2}(l_{\ast_{1}}(x)b-r_{\ast_{1}}(x)b)+r_{\prec_{1}}(r_{\ast_{2}}(b)x-l_{\ast_{2}}(b)x)a=l_{\ast_{1}}(x)(a\prec_{2}b)-(l_{\ast_{1}}(x)a)\prec_{2}b-l_{\prec_{1}}(r_{\ast_{2}}(a)x)b,$
    (42) $l_{\prec_{1}}(x)(a\ast_{2}b-b\ast_{2}a)=a\ast_{2}(l_{\prec_{1}}(x)b)+r_{\ast_{1}}(r_{\prec_{2}}(b)x)a-(r_{\ast_{1}}(x)a)\prec_{2}b-l_{\prec_{1}}(l_{\ast_{2}}(a)x)b,$
    (43) $r_{\ast_{1}}(x)(a\cdot_{2}b)=(r_{\ast_{1}}(x)a)\succ_{2}b+l_{\prec_{1}}(l_{\ast_{2}}(a)x)b+a\succ_{2}(r_{\ast_{1}}(x)b)+r_{\succ_{1}}(l_{\ast_{2}}(b)x)a,$
    (44) $(r_{\cdot_{1}}(x)a)\ast_{2}b+l_{\ast_{1}}(l_{\cdot_{2}}(a)x)b=r_{\prec_{1}}(x)(a\ast_{2}b)+a\succ_{2}(l_{\ast_{1}}(x)b)+r_{\succ_{1}}(r_{\ast_{2}}(b)x)a,$
    (45) $(l_{\cdot_{1}}(x)a)\ast_{2}b+l_{\ast_{1}}(r_{\cdot_{2}}(a)x)b=(l_{\ast_{1}}(x)b)\prec_{2}a+l_{\prec_{1}}(r_{\ast_{2}}(b)x)a+l_{\succ_{1}}(x)(a\ast_{2}b),$
where $\cdot_{1}=\succ_{1}+\prec_{1},~\cdot_{2}=\succ_{2}+\prec_{2},~[x,y]=x\ast_{1}y-y\ast_{1}x$ and $a,b]=a\ast_{2}b-b\ast_{2}a$. In this case, we denote this noncommutative pre-Poisson algebra by $A_{1}\bowtie A_{2}$, and $(A_{1},A_{2},l_{\succ_{1}},r_{\succ_{1}},l_{\prec_{1}},r_{\prec_{1}},l_{\ast_{1}},r_{\ast_{1}},$ $l_{\succ_{2}},r_{\succ_{2}},l_{\prec_{2}},r_{\prec_{2}},l_{\ast_{2}},r_{\ast_{2}})$ satisfying the above conditions is called a matched pair of noncommutative pre-Poisson algebras.
Proof.
To prove that $(A_{1}\oplus A_{2},\succ,\prec,\ast)$ is a noncommutative pre-Poisson algebra, we only need to check that Eqs. (3)-(5) hold for all $x+a,y+b,z+c\in A_{1}\oplus A_{2}$ with $x,y,z\in A_{1},a,b,c\in A_{2}$. Indeed, we may verify the following cases through direct computations:
  1. (a)
    Eq. (3) holds for $(x,y,a)$ if and only if Eqs. (7) and (28) hold.
  2. (b)
    Eq. (3) holds for $(x,a,y)$ if and only if Eqs. (8) and (29) hold.
  3. (c)
    Eq. (3) holds for $(a,x,y)$ if and only if Eqs. (13) and (30) hold.
  4. (d)
    Eq. (3) holds for $(a,b,x)$ if and only if Eqs. (7) and (37) hold.
  5. (e)
    Eq. (3) holds for $(a,x,b)$ if and only if Eqs. (8) and (38) hold.
  6. (f)
    Eq. (3) holds for $(x,a,b)$ if and only if Eqs. (13) and (39) hold.
  7. (g)
    Eq. (4) holds for $(x,y,a)$ if and only if Eqs. (9) and (31) hold.
  8. (h)
    Eq. (4) holds for $(x,a,y)$ if and only if Eqs. (14) and (32) hold.
  9. (i)
    Eq. (4) holds for $(a,x,y)$ if and only if Eqs. (10) and (33) hold.
  10. (j)
    Eq. (4) holds for $(a,b,x)$ if and only if Eqs. (9) and (40) hold.
  11. (k)
    Eq. (4) holds for $(a,x,b)$ if and only if Eqs. (14) and (41) hold.
  12. (l)
    Eq. (4) holds for $(x,a,b)$ if and only if Eqs. (10) and (42) hold.
  13. (m)
    Eq. (5) holds for $(x,y,a)$ if and only if Eqs. (15) and (34) hold.
  14. (n)
    Eq. (5) holds for $(x,a,y)$ if and only if Eqs. (11) and (35) hold.
  15. (o)
    Eq. (5) holds for $(a,x,y)$ if and only if Eqs. (12) and (36) hold.
  16. (p)
    Eq. (5) holds for $(a,b,x)$ if and only if Eqs. (15) and (43) hold.
  17. (q)
    Eq. (5) holds for $(a,x,b)$ if and only if Eqs. (11) and (44) hold.
  18. (r)
    Eq. (5) holds for $(x,a,b)$ if and only if Eqs. (12) and (45) hold.
Remark 2.11.
  1. (a)
    If the linear maps $l_{\succ_{2}},r_{\succ_{2}},l_{\prec_{2}},r_{\prec_{2}},l_{\ast_{2}},r_{\ast_{2}}:A_{2}\rightarrow\text{End}(A_{1})$ are trivial, then the matched pair reduces to an A-noncommutative pre-Poisson algebra. For the completeness, we provide the details. Let $(A_{1},\succ_{1},\prec_{1},\ast_{1})$ and $(A_{2},\succ_{2},\prec_{2},\ast_{2})$ be two noncommutative pre-Poisson algebras. Assume that there are linear maps
    $l_{\succ_{1}},r_{\succ_{1}},l_{\prec_{1}},r_{\prec_{1}},l_{\ast_{1}},r_{\ast_{1}}:A_{1}\longrightarrow\hbox{End}(A_{2})$
    such that $(A_{2},l_{\succ_{1}},r_{\succ_{1}},l_{\prec_{1}},r_{\prec_{1}},l_{\ast_{1}},r_{\ast_{1}})$ is a representation of $(A_{1},\succ_{1},\prec_{1},\ast_{1})$, and the following compatible conditions hold for all $x,y\in A_{1},a,b\in A_{2}$:
    (46) $r_{\succ_{1}}(x)(a\ast_{2}b-b\ast_{2}a)=a\ast_{2}r_{\succ_{1}}(x)b-b\succ_{2}(r_{\ast_{1}}(x)a),$
    (47) $(r_{\ast_{1}}(x)a-l_{\ast_{1}}(x)a)\succ_{2}b=a\ast_{2}(l_{\succ_{1}}(x)b)-l_{\succ_{1}}(x)(a\ast_{2}b),$
    (48) $(l_{\ast_{1}}(x)a-r_{\ast_{1}}(x)a)\succ_{2}b=l_{\ast_{1}}(x)(a\succ_{2}b)-a\succ_{2}(l_{\ast_{1}}(x)b),$
    (49) $a\prec_{2}(r_{\ast_{1}}(x)b-l_{\ast_{1}}(x)b)=b\ast_{2}(r_{\prec_{1}}(x)a)-r_{\prec_{1}}(x)(b\ast_{1}a),$
    (50) $a\prec_{2}(l_{\ast_{1}}(x)b-r_{\ast_{1}}(x)b)=l_{\ast_{1}}(x)(a\prec_{2}b)-(l_{\ast_{1}}(x)a)\prec_{2}b,$
    (51) $l_{\prec_{1}}(x)(a\ast_{2}b-b\ast_{2}a)=a\ast_{2}(l_{\prec_{1}}(x)b)-(r_{\ast_{1}}(x)a)\prec_{2}b,$
    (52) $r_{\ast_{1}}(x)(a\cdot_{2}b)=(r_{\ast_{1}}(x)a)\succ_{2}b+a\succ_{2}(r_{\ast_{1}}(x)b),$
    (53) $(r_{\cdot_{1}}(x)a)\ast_{2}b=r_{\prec_{1}}(x)(a\ast_{2}b)+a\succ_{2}(l_{\ast_{1}}(x)b),$
    (54) $(l_{\cdot_{1}}(x)a)\ast_{2}b=(l_{\ast_{1}}(x)b)\prec_{2}a+l_{\succ_{1}}(x)(a\ast_{2}b),$
    (55) $r_{\ast_{1}}(x)(a\ast_{2}b-b\ast_{2}a)=a\ast_{2}(r_{\ast_{1}}(x)b)-b\ast_{2}(r_{\ast_{1}}(x)a),$
    (56) $l_{\ast_{1}}(x)(a\ast_{2}b)=(l_{\ast_{1}}(x)-r_{\ast_{1}}(x)a)\ast_{2}b+a\ast_{2}(l_{\ast_{1}}(x)b),$
    (57) $(l_{\prec_{1}}(x)a)\prec_{2}b=l_{\prec_{1}}(x)(a\cdot_{2}b),\ \ \ (l_{\succ_{1}}(x)a)\prec_{2}b=l_{\succ_{1}}(x)(a\prec_{2}b),\ \ \$
    (58) $l_{\succ_{1}}(x)(a\succ_{2}b)=(l_{\cdot_{1}}(x)a)\succ_{2}b,\ \ \ (r_{\prec_{1}}(x)a)\prec_{2}b=a\prec_{2}(l_{\cdot_{1}}(x)b),$
    (59) $(r_{\succ_{1}}(x)a)\prec_{2}b=a\succ_{2}(l_{\prec_{1}}(x)b),\ \ \ a\succ_{2}l_{\succ}(x)b=r_{\cdot}(x)a\succ_{2}b,$
    (60) $r_{\prec_{1}}(x)(a\prec_{2}b)=a\prec_{2}(r_{\cdot_{1}}(x)b),\ \ \ r_{\prec_{1}}(x)(a\succ_{2}b)=a\succ_{2}(r_{\prec_{1}}(x)b),$
    (61) $a\succ_{2}(r_{\succ_{1}}(x)b)=r_{\succ_{1}}(x)(a\cdot_{2}b).$
    Then $(A_{2},l_{\succ_{1}},r_{\succ_{1}},l_{\prec_{1}},r_{\prec_{1}},l_{\ast_{1}},r_{\ast_{1}})$ is called an A-noncommutative pre-Poisson algebra. If only Eqs. (55)-(56) are satisfied, $(A_{2},l_{\ast_{1}},r_{\ast_{1}})$ is called an A-pre-Lie algebra. If only Eqs. (57)-(61) are satisfied, $(A_{2},l_{\succ_{1}},r_{\succ_{1}},l_{\prec_{1}},r_{\prec_{1}})$ is called an A-dendriform algebra. A-dendriform algebra is named an action of $A_{1}$ on $A_{2}$ in [30].
  2. (b)
    If the linear maps $l_{\succ_{2}},r_{\succ_{2}},l_{\prec_{2}},r_{\prec_{2}},l_{\ast_{2}},r_{\ast_{2}}:A_{2}\rightarrow\text{End}(A_{1})$ and $(\succ_{2},\prec_{2},\ast_{2})$ are trivial, then the matched pair reduces to the semidirect product. Denote it simply by $A_{1}\ltimes A_{2}$. In particular, if in addition, $A_{1}$ is coherent and $(l_{\succ_{1}},r_{\succ_{1}},l_{\prec_{1}},r_{\prec_{1}},l_{\ast_{1}},r_{\ast_{1}})$ satisfying Eq. (18)-(20), then $A_{1}\ltimes A_{2}$ and $A_{1}\ltimes A_{2}^{*}$ both are coherent noncommutative pre-Poisson algebras.
2.2. Phase spaces of noncommutative Poisson algebras
We now introduce the notions of symplectic noncommutative Poisson algebras and the phase spaces of noncommutative Poisson algebras.
Definition 2.12.
A symplectic noncommutative Poisson algebra is a noncommutative Poisson algebra $(A,\cdot,[\ ,\ ])$ together with a non-degenerate skew-symmetric bilinear form $\omega$ satisfying the following conditions for all $x,y,z\in A$,
(62) $\omega([x,y],z)+\omega([y,z],x)+\omega([z,x],y)=0,$
(63) $\omega(x\cdot y,z)+\omega(y\cdot z,x)+\omega(z\cdot x,y)=0,$
that is, $(A,[\ ,\ ],\omega)$ is a symplectic Lie algebra [3] and $(A,\cdot,\omega)$ is an associative algebra with a Connes cocycle [4].
Definition 2.13.
A quadratic noncommutative pre-Poisson algebra is a noncommutative pre-Poisson algebra $(A,\succ,\prec,\ast)$ endowed with a nondegenerate skew-symmetric bilinear form $\omega:A\otimes A\longrightarrow k$ satisfying the following invariance conditions for all $x,y,z\in A$
(64) $\omega(x\ast y,z)=-\omega(y,[x,z]),$
(65) $\omega(x\succ y,z)=\omega(y,z\cdot x),\ \ \ \omega(x\prec y,z)=\omega(x,y\cdot z),$
where $x\ast z-z\ast x=[x,z]$ and $x\cdot y=x\succ y+x\prec y$.
Example 2.14.
Let $(A,\succ,\prec,\ast)$ be the 2-dimensional noncommutative pre-Poisson algebra with a basis $\{e_{1},e_{2}\}$ given in Example 2.3. Define a nondegenerate skew-symmetric bilinear form $\omega$ on $A$ as follows: $\omega(e_{1},e_{2})=-\omega(e_{2},e_{1})=1$ and $\omega(e_{2},e_{2})=\omega(e_{1},e_{1})=0$. Then $(A,\succ,\prec,\ast,\omega)$ is a quadratic noncommutative pre-Poisson algebra.
Proposition 2.15.
Let $(A,\cdot,[\ ,\ ],\omega)$ be a symplectic noncommutative Poisson algebra. Then there is a compatible coherent noncommutative pre-Poisson algebra $(A,\succ,\prec,\ast)$ given by equations (64)-(65). Conversely, assume that $(A,\succ,\prec,\ast,\omega)$ is a quadratic noncommutative pre-Poisson algebra and $(A,\cdot,[\ ,\ ])$ is its sub-adjacent noncommutative Poisson algebra. Then $(A,\cdot,[\ ,\ ],\omega)$ is a symplectic noncommutative Poisson algebra.
Proof.
By Theorem 4.1.1 [4] and Theorem 2.2 [3], we get that $(A,\succ,\prec)$ is a dendriform algebra and $(A,\ast)$ is a pre-Lie algebra. Using Eqs. (1) and (64)-(65), we have for all $x,y,z,w\in A$,
$\omega((x\ast y-y\ast x)\succ z,w)-\omega(x\ast(y\succ z)-y\succ(x\ast z),w)$
$=$ $\omega(z,w\cdot[x,y])+\omega(z,[x,w]\cdot y)-\omega(z,[x,w\cdot y])$
$=$ $0.$
Thus, Eq. (3) holds. Analogously, Eqs. (4)-(5) hold. By Eqs. (64)-(65) and (1), we get
$\omega((x\cdot y)\ast z-x\ast(y\succ z)-y\ast(z\prec x),w)$
$=$ $-\omega(z,[x\cdot y,w])+\omega(y\succ z,[x,w])+\omega(z\prec x,[y,w])$
$=$ $-\omega(z,[x\cdot y,w])+\omega(z,[x,w]\cdot y)+\omega(z,x\cdot[y,w])$
$=$ $0,$
which yields that Eq. (6) holds, that is, $(A,\succ,\prec,\ast)$ is coherent. The converse part is easy to check. The proof is completed. ∎
Definition 2.16.
Let $(A,\cdot_{A},[\ ,\ ]_{A})$ be a noncommutative Poisson algebra and $A^{*}$ its dual space. If there is a noncommutative Poisson algebra structure $(\cdot,[\ ,\ ])$ on the direct sum vector space $A\oplus A^{*}$ such that $(A\oplus A^{*},\cdot,[\ ,\ ],\omega)$ is a symplectic noncommutative Poisson algebra, where $\omega$ is given by the following equation:
(66) $\omega(x+a,y+b)=\langle a,y\rangle-\langle x,b\rangle,\forall~x,y\in A,~a,b\in A^{*},$
and both $(A,\cdot_{A},[\ ,\ ]_{A})$, $(A^{*},\cdot|_{A^{*}},[\ ,\ ]|_{A^{*}})$ are noncommutative Poisson subalgebras of $(A\oplus A^{*},\cdot,[\ ,\ ])$, then the symplectic noncommutative Poisson algebra $(A\oplus A^{*},\cdot,[\ ,\ ],\omega)$ is called a phase space of the noncommutative Poisson algebra $(A,\cdot_{A},[\ ,\ ]_{A})$.
Proposition 2.17.
A noncommutative Poisson algebra has a phase space if and only if there is a compatible coherent noncommutative pre-Poisson algebra.
Proof.
Assume that $(A\oplus A^{*},\cdot,[\ ,\ ],\omega)$ is a phase space of the noncommutative Poisson algebra $(A,\cdot_{A},[\ ,\ ]_{A})$. In light of Proposition 2.15, there is a compatible coherent noncommutative pre-Poisson algebra structure $(\succ,\prec,\ast)$ on $A\oplus A^{*}$ defined by Eqs. (64)-(65). Due to $(A,\cdot_{A},[\ ,\ ]_{A})$ being a noncommutative Poisson subalgebra of $A\oplus A^{*}$, we obtain for all $x,y,z\in A$,
$\omega(x\ast y,z)=-\omega(y,[x,z])=0,$
$\omega(x\succ y,z)=\omega(y,z\cdot x)=0,\ \ \ \omega(x\prec y,z)=\omega(x,y\cdot z)=0,$
which means that $x\ast y,x\succ y,x\prec y\in A$, which indicates that $(A,\succ|_{A},\prec|_{A},\ast|_{A})$ and $(A,\succ_{A},\prec_{A},\ast_{A})$ are consistent. Thus, the sub-adjacent noncommutative Poisson algebra of $(A,\succ|_{A},\prec|_{A},\ast|_{A})$ is exactly $(A,\cdot_{A},[\ ,\ ]_{A})$.
Conversely, assume that $(A,\succ_{A},\prec_{A},\ast_{A})$ is a coherent noncommutative pre-Poisson algebra. By Example 2.7, we know that $(A^{*},R_{\prec}^{*},L_{\succ}^{*},-L_{\ast}^{*})$ is a representation of $(A,\cdot_{A},[\ ,\ ]_{A})$. It follows that $(A\ltimes A^{*},\cdot,[\ ,\ ])$ is a noncommutative Poisson algebra, where
$(x+a)\cdot(y+b)=x\succ_{A}y+R_{\prec}^{*}(x)b+L_{\succ}^{*}(y)a,$
$[x+a,y+b]=[x,y]_{A}-L_{\ast}^{*}(x)b+L_{\ast}^{*}(y)a,$
for all $x,y\in A,a,b\in A^{*}$. It is easy to check that $(A\ltimes A^{*},\cdot,[\ ,\ ],\omega)$ is a a phase space of the noncommutative Poisson algebra $(A,\cdot_{A},[\ ,\ ]_{A})$, where $\omega$ is defined by Eq. (66). ∎
Corollary 2.18.
Let $(A\oplus A^{*},\cdot,[\ ,\ ],\omega)$ be a phase space of the noncommutative Poisson algebra $(A,\cdot_{A},[\ ,\ ]_{A})$ and $(A\oplus A^{*},\succ,\prec,\ast)$ the compatible noncommutative pre-Poisson algebra. Then both $(A,\succ|_{A},\prec|_{A},\ast|_{A})$ and $(A^{*},\succ|_{A^{*}},\prec|_{A^{*}},\ast|_{A^{*}})$ are subalgebras of the noncommutative pre-Poisson algebra $(A\oplus A^{*},\succ,\prec,\ast)$.
Proof.
This follows directly from the proof of Proposition 2.17. ∎
Definition 2.19.
A Manin triple of noncommutative pre-Poisson algebras is a quadruple $(A,A_{1},A_{2},\omega)$, where $(A,\omega)$ is a quadratic noncommutative pre-Poisson algebra, $(A_{1},\succ_{1},\prec_{1},\ast_{1})$ and $(A_{2},\succ_{2},\prec_{2},\ast_{2})$ are noncommutative pre-Poisson subalgebras of $A$, such that
  1. (a)
    $A=A_{1}\oplus A_{2}$ as vector spaces;
  2. (b)
    $A_{1}$ and $A_{2}$ are isotropic with respect to $\omega$.
Proposition 2.20.
Let $(A,A_{1},A_{2},\omega)$ be a Manin triple of noncommutative pre-Poisson algebras. Then the noncommutative pre-Poisson algebras $A$, $A_{1}$ and $A_{2}$ are coherent.
Proof.
By Eqs. (64)-(65) and (3)-(5), we have for all $x,y,z\in A_{1}$ and $a\in A_{2}$,
$\omega((x\cdot y)\ast z,a)-\omega(x\ast(y\succ z),a)-\omega(y\ast(z\prec x),a)$
$=$ $-\omega(z,(x\cdot y)\ast a-a\ast(x\cdot y))+\omega(y\succ z,x\ast a-a\ast x)+\omega(z\prec x,y\ast a-a\ast y)$
$=$ $-\omega(z,(x\cdot y)\ast a-a\ast(x\cdot y))+\omega(z,(x\ast a-a\ast x)\cdot y)+\omega(z,x\cdot(y\ast a-a\ast y))$
$=$ $0,$
which indicates that Eq. (6) holds. Thus, $A_{1}$ is coherent. Analogously, $A$ and $A_{2}$ are coherent. ∎
Theorem 2.21.
Suppose that $(A,\succ_{A},\prec_{A},\ast_{A})$ and $(A^{*},\succ_{A^{*}},\prec_{A^{*}},\ast_{A^{*}})$ are two noncommutative pre-Poisson algebras and their sub-adjacent noncommutative Poisson algebras are $(A,\cdot_{A},[\ ,\ ]_{A})$ and $(A^{*},\cdot_{A^{*}},[\ ,\ ]_{A^{*}})$ respectively. Then the following conditions are equivalent:
  1. (a)
    $(A\oplus A^{*},\cdot,[\ ,\ ],\omega)$ is a phase space of the noncommutative Poisson algebra $(A,\cdot_{A},[\ ,\ ]_{A})$.
  2. (b)
    $(A\oplus A^{*},A,A^{*},\omega)$ is a Manin triple of noncommutative pre-Poisson algebras with the bilinear form given by Eq. (66) and the isotropic subalgebras are $A$ and $A^{*}$.
  3. (c)
    $(A,A^{*},R_{\cdot_{1}}^{*},-L_{\prec_{1}}^{*},-R_{\succ_{1}}^{*},L_{\cdot_{1}}^{*},-\mathrm{ad_{1}}^{*},R^{*}_{\ast_{1}},R_{\cdot_{2}}^{*},-L_{\prec_{2}}^{*},-R_{\succ_{2}}^{*},L_{\cdot_{2}}^{*},-\mathrm{ad_{2}}^{*},R^{*}_{\ast_{2}})$ is a matched pair of coherent noncommutative pre-Poisson algebras.
  4. (d)
    $(A,A^{*},R_{\prec_{1}}^{*},L_{\succ_{1}}^{*},-L_{\ast_{1}}^{*},R_{\prec_{2}}^{*},L_{\succ_{2}}^{*},-L_{\ast_{2}}^{*})$ is a matched pair of coherent noncommutative Poisson algebras.
Proof.
(a) $\Longrightarrow$ (b) Assume that $(A\oplus A^{*},\cdot,[\ ,\ ],\omega)$ is a phase space of the noncommutative Poisson algebra $(A,\cdot_{A},[\ ,\ ]_{A})$. According to Proposition 2.15, there is a coherent noncommutative pre-Poisson algebra structure $(\succ,\prec,\ast)$ on $A\oplus A^{*}$ defined by Eqs. (64)-(65) such that $(A\oplus A^{*},\succ,\prec,\ast)$ is quadratic noncommutative pre-Poisson algebra. In view of Corollary 2.18, $(A,\succ|_{A},\prec|_{A},\ast|_{A})$ and $(A^{*},\succ|_{A^{*}},\prec|_{A^{*}},\ast|_{A^{*}})$ are subalgebras of $(A\oplus A^{*},\succ,\prec,\ast)$. It is obvious that $A$ and $A^{*}$ are isotropic. Thus, $(A\oplus A^{*},A,A^{*},\omega)$ is a Manin triple of noncommutative pre-Poisson algebras. Combining Proposition 4.2 [3], Theorem 4.1.5 [4] and Proposition 2.10, it follows directly that (b) $\Longleftrightarrow$ (c) $\Longleftrightarrow$ (d) $\Longrightarrow$ (a). ∎
3. Noncommutative pre-Poisson bialgebras and the noncommutative pre-Poisson Yang-Baxter equation
3.1. Dendriform bialgebras and pre-Lie bialgebras
We start by reviewing the bialgebra theories for dendriform algebras (see [4]) and pre-Lie bialgebras (see [3]).
Definition 3.1.
  1. (a)
    A dendriform coalgebra is a triple $(A,\Delta_{\succ},\Delta_{\prec})$, where $A$ is a vector space and $\Delta_{\succ},\Delta_{\prec}:A\longrightarrow A\otimes A$ are linear maps such that the following conditions hold:
    $(\Delta_{\succ}\otimes I)\Delta_{\prec}=(I\otimes\Delta_{\prec})\Delta_{\succ},\ \ \ (I\otimes\Delta_{\succ})\Delta_{\succ}=(\Delta\otimes I)\Delta_{\succ},\ \ \ (\Delta_{\prec}\otimes I)\Delta_{\prec}=(I\otimes\Delta)\Delta_{\prec},$
    where $\Delta=\Delta_{\succ}+\Delta_{\prec}.$
  2. (b)
    A dendriform bialgebra is a quintuple $(A,\succ,\prec,\Delta_{\succ},\Delta_{\prec})$, where $(A,\succ,\prec)$ is a dendriform algebra and $(A,\Delta_{\succ},\Delta_{\prec})$ is a dendriform coalgebra such that for all $x,y\in A$, the following compatible conditions hold:
    $\Delta_{\prec}(x\cdot_{A}y)=((I\otimes L_{\succ_{A}}(x))\Delta_{\prec_{A}}(y)+R_{\cdot_{A}}(y)\otimes I)\Delta_{\prec}(x),$
    $\Delta_{\succ}(x\cdot_{A}y)=(I\otimes L_{\cdot_{A}}(x))\Delta_{\succ}(y)+(R_{\prec_{A}}(y)\otimes I)\Delta_{\succ}(x),$
    $\Delta(x\prec y)=(R_{\prec}(y)\otimes I)\Delta(x)+(I\otimes L_{\prec}(x))\Delta_{\succ}(y),$
    $\Delta(x\succ y)=(I\otimes L_{\succ}(x))\Delta(y)+(R_{\succ}(y)\otimes I)\Delta_{\prec}(x),$
    $(L_{\cdot_{A}}(x)\otimes I-(I\otimes R_{\prec_{A}}(x))\Delta_{\prec}(y)+\tau(L_{\succ_{A}}(y)\otimes I-I\otimes R_{\cdot_{A}}(y))\Delta_{\succ}(x)=0,$
    $(L_{\succ}(x)\otimes I)\Delta(y)-(R_{\succ}(y)\otimes I)\tau\Delta_{\succ}(x)+(I\otimes L_{\prec}(y))\tau\Delta_{\prec}(x)-(I\otimes R_{\prec}(x))\Delta(y)=0,$
    where $\cdot=\succ+\prec,~\Delta=\Delta_{\succ}+\Delta_{\prec}$ and $R_{\cdot}=R_{\prec}+R_{\succ},~L_{\cdot}=L_{\prec}+L_{\succ}$.
Theorem 3.2.
Let $(A,\succ,\prec)$ be a dendriform algebra and $r=\sum_{i}a_{i}\otimes b_{i}\in A\otimes A$. Assume that $\Delta_{\succ,r},\Delta_{\prec,r}$ defined by
(67) $\Delta_{\succ,r}(x)=(I\otimes L_{\cdot}(x)-R_{\prec}(x)\otimes I)\tau(r),$
(68) $\Delta_{\prec,r}(x)=(R_{\cdot}(x)\otimes I-I\otimes L_{\succ}(x))r,$
Then $(A,\succ,\prec,\Delta_{\succ,r},\Delta_{\prec,r})$ is a dendriform bialgebra if and only if the following equations hold:
(69) $[Q(x\succ y)-I\otimes L_{\succ}(x)Q(y)](r-\tau(r))=0,$
(70) $\tau(Q(x))Q(y)(r-\tau(r))=0,$
(71) $(R_{\cdot}(x)\otimes I\otimes I-I\otimes I\otimes L_{\succ}(x))(r_{12}\cdot r_{13}-r_{23}\succ r_{12}-r_{13}\prec r_{32})$
$+\sum_{i}(a_{i}\cdot x)\otimes Q(a_{i})(r-\tau(r))-a_{i}\otimes[Q(x\succ b_{i})(r-\tau(r))]=0,$
(72) $(R_{\prec}(x)\otimes I\otimes I-I\otimes I\otimes L_{\succ}(x))(r_{31}\succ r_{23}+r_{21}\prec r_{13}-r_{23}\cdot r_{21})=0,$
(73) $(R_{\prec}(x)\otimes I\otimes I-I\otimes I\otimes L_{\cdot}(x))(r_{32}\prec r_{21}+r_{12}\succ r_{31}-r_{31}\cdot r_{32})$
$+[Q(b_{i})(r-\tau(r))\otimes x\cdot a_{i}-Q(b_{i}\prec x)(r-\tau(r))\otimes a_{i}]=0,$
where $Q(x)=I\otimes L_{\succ}(x)-R_{\prec}(x)\otimes I$.
Proposition 3.3.
Let $(A,\succ,\prec)$ be a dendriform algebra and $r=\sum\limits_{i}a_{i}\otimes b_{i}\in A\otimes A$ symmetric. Define linear maps $\Delta_{\succ,r},\Delta_{\prec,r}:A\longrightarrow A\otimes A$ by Eqs. (67)-(68). Then $(A,\succ,\prec,\ast,\Delta_{\succ,r},\Delta_{\prec,r})$ is a dendriform bialgebra if and only if the following equations hold:
(74) $(R_{\cdot}(x)\otimes I\otimes I-I\otimes I\otimes L_{\succ}(x))(r_{12}\cdot r_{13}-r_{23}\succ r_{12}-r_{13}\prec r_{23})=0,$
(75) $(R_{\prec}(x)\otimes I\otimes I-I\otimes I\otimes L_{\succ}(x))(r_{13}\succ r_{23}+r_{12}\prec r_{13}-r_{23}\cdot r_{12})=0,$
(76) $(R_{\prec}(x)\otimes I\otimes I-I\otimes I\otimes L_{\cdot}(x))(r_{23}\prec r_{12}+r_{12}\succ r_{13}-r_{13}\cdot r_{23})=0.$
Let $(A,\succ,\prec)$ be a dendriform algebra and $r\in A\otimes A$. The equation
(77) $D(r)=r_{12}\cdot r_{13}-r_{23}\succ r_{12}-r_{13}\prec r_{23}=0$
is called the D-equation in $(A,\succ,\prec)$.
Theorem 3.4.
Let $(A,\succ,\prec)$ be a dendriform algebra and $r\in A\otimes A$. If $r$ is a symmetric solution of the D-equation in $(A,\succ,\prec)$, then $(A,\succ,\prec,\Delta_{\succ,r},\Delta_{\prec,r})$ is a dendriform bialgebra.
Definition 3.5.
  1. (a)
    A pre-Lie coalgebra is a vector space $A$ together with a linear map $\delta:A\longrightarrow A\otimes A$ such that
    $(I\otimes\delta)\delta-(\tau\otimes I)(I\otimes\delta)\delta=(\delta\otimes I)\delta-(\tau\otimes I)(\delta\otimes I)\delta.$
  2. (b)
    A pre-Lie bialgebra is a triple $(A,\ast,\delta)$, where $(A,\ast)$ is a pre-Lie algebra and $(A,\delta)$ is a pre-Lie coalgebra such that the following compatible conditions hold for all $x,y\in A$,
    $\delta(x\ast y-y\ast x)=(I\otimes\mathrm{ad}(x)+L_{\ast}(x)\otimes I)\delta(y)-(I\otimes\mathrm{ad}(y)+L_{\ast}(y)\otimes I)\delta(x),$
    $(\delta-\tau\delta)(x\ast y)=L_{\ast}(x)\otimes I+I\otimes L_{\ast}(x))(\delta-\tau\delta)(y)+(I\otimes L_{\ast}(y))\delta(x)-(L_{\ast}(y)\otimes I)\tau\delta(x).$
Theorem 3.6.
Let $(A,\ast)$ be a pre-Lie algebra and $r=\sum_{i}a_{i}\otimes b_{i}\in A\otimes A$. Define a linear map $\delta_{r}:A\rightarrow A\otimes A$ by
(78) $\delta_{r}(x)=-(L_{\ast}(x)\otimes I+I\otimes\mathrm{ad}(x))r,\;\forall x\in A.$
Then $(A,\ast,\delta_{r})$ is a pre-Lie bialgebra if and only if the following equations hold:
(79) $(L_{\ast}(x)\otimes I\otimes I+I\otimes L_{\ast}(x)\otimes I+I\otimes\otimes I\otimes\mathrm{ad}(x))[[r,r]]=0,$
(80) $(P(x\ast y)-P(x)P(y)(r-\tau(r))=0,$
where $P(x)=L_{\ast}(x)\otimes I+I\otimes L_{\ast}(x)$ and $r,r]]=r_{13}\ast r_{12}+[r_{23},r_{12}]-r_{23}\ast r_{21}-[r_{13},r_{21}]+[r_{23},r_{13}]$.
Proposition 3.7.
Let $(A,\ast)$ be a pre-Lie algebra and $r=\sum\limits_{i}a_{i}\otimes b_{i}\in A\otimes A$ symmetric. Assume that $\delta_{r}:A\longrightarrow A\otimes A$ is given by Eq. (78). Then $(A,\ast,\delta_{r})$ is a pre-Lie bialgebra if and only if the following equations hold:
(81) $(L_{\ast}(x)\otimes I\otimes I+I\otimes L_{\ast}(x)\otimes I+I\otimes\otimes I\otimes\mathrm{ad}(x))S(r)=0,$
where
$S(r)=r_{12}\ast r_{23}-r_{12}\ast r_{13}+[r_{13},r_{23}]=0.$
Let $(A,\ast)$ be a pre-Lie algebra and $r\in A\otimes A$. The equation $S(r)=0$ is called the S-equation in $(A,\ast)$.
Theorem 3.8.
Let $(A,\ast)$ be a pre-Lie algebra and $r\in A\otimes A$. If $r$ is a symmetric solution of the S-equation in $(A,\ast)$, then $(A,\ast,\delta_{r})$ is a pre-Lie bialgebra.
3.2. Noncommutative pre-Poisson bialgebras
This section develops the bialgebra theory for noncommutative pre-Poisson algebras.
Definition 3.9.
A noncommutative pre-Poisson coalgebra is a quadruple $(A,\Delta_{\succ},\Delta_{\prec},\delta)$, where $(A,\Delta_{\succ},\Delta_{\prec})$ is a dendriform coalgebra and $(A,\delta)$ is a pre-Lie coalgebra and $\Delta_{\succ},\Delta_{\prec},\delta$ are compatible in the following sense:
(82) $(\delta\otimes I)\Delta_{\succ}-(\tau\otimes I)(\delta\otimes I)\Delta_{\succ}=(I\otimes\Delta_{\succ})\delta-(\tau\otimes I)(I\otimes\delta)\Delta_{\succ},$
(83) $(I\otimes\delta)\Delta_{\prec}-(I\otimes\tau)(I\otimes\delta)\Delta_{\prec}=(\tau\otimes I)(I\otimes\Delta_{\prec})\delta-(\tau\otimes I)(\delta\otimes I)\Delta_{\prec}$
(84) $(\Delta_{\succ}\otimes I)\delta+(\Delta_{\prec}\otimes I)\delta=(I\otimes\tau)(\delta\otimes I)\Delta_{\prec}+(I\otimes\delta)\Delta_{\succ}.$
Definition 3.10.
A noncommutative pre-Poisson coalgebra $(A,\Delta_{\succ},\Delta_{\prec},\delta)$ is called coherent if it satisfies
(85) $(\Delta_{\succ}\otimes I+\Delta_{\prec}\otimes I)\delta=(I\otimes\Delta_{\succ})\delta+(\tau\otimes I)(I\otimes\tau)(I\otimes\Delta_{\prec})\delta.$
Definition 3.11.
Let $(A,\succ,\prec,\ast)$ be a coherent noncommutative pre-Poisson algebra. Suppose that there are comultiplications $\Delta_{\succ},\Delta_{\prec},\delta:A\longrightarrow A\otimes A$ such that $(A,\Delta_{\succ},\Delta_{\prec},\delta)$ is a coherent noncommutative pre-Poisson coalgebra. If in addition, $(A,\succ,\prec,\Delta_{\succ},\Delta_{\prec})$ is a dendriform bialgebra, $(A,\ast,\delta)$ is a pre-Lie bialgebra and $\Delta_{\succ},\Delta_{\prec},\delta$ satisfy the following compatible conditions:
(86) $\Delta(x\ast y)=(L_{\ast}(x)\otimes I)\Delta(y)+(I\otimes L_{\ast}(x))\Delta(y)-(I\otimes L_{\prec}(y))\delta(x)-(R_{\succ}(y)\otimes I)\tau\delta(x),$
(87) $(\delta-\tau\delta)(y\prec x)=(I\otimes R_{\prec}(x))(\delta-\tau\delta)(y)+(R_{\ast}(y)\otimes I)\tau\Delta_{\succ}(x)+(I\otimes L_{\prec}(y))\delta(x)-(L_{\ast}(x)\otimes I)\Delta(y),$
(88) $\delta(x\cdot y)=(I\otimes R_{\cdot}(y))\delta(x)+(I\otimes L_{\cdot}(x))\delta(y)-(L_{\ast}(x)\otimes I)\tau\Delta_{\prec}(y)-(L_{\ast}(y)\otimes I)\Delta_{\succ}(x),$
(89) $\Delta_{\prec}(x\ast y-y\ast x)=(\mathrm{ad}(x)\otimes I)\Delta_{\prec}(y)+(I\otimes L_{\succ}(y))\tau\delta(x)+(I\otimes L_{\ast}(x))\Delta_{\prec}(y)-(R_{\cdot}(y)\otimes I)\tau\delta(x),$
for all $x,y\in A$, then $(A,\succ,\prec,\ast,\Delta_{\succ},\Delta_{\prec},\delta)$ is called a noncommutative pre-Poisson bialgebra.
Remark 3.12.
$(A,\Delta_{\succ},\Delta_{\prec},\delta)$ is a (coherent) noncommutative pre-Poisson coalgebra if and only if $(A^{*},\succ_{A^{*}},\prec_{A^{*}},\ast_{A^{*}})$ is a (coherent) noncommutative pre-Poisson algebra, where $\succ_{A^{*}},\prec_{A^{*}},\ast_{A^{*}}$ are the linear dual of $\Delta_{\succ},\Delta_{\prec},\delta$ respectively, that is,
(90) $\langle\Delta_{\succ}(x),\zeta\otimes\eta\rangle=\langle x,\zeta\succ_{A^{*}}\eta\rangle,\ \ \ \langle\Delta_{\prec}(x),\zeta\otimes\eta\rangle=\langle x,\zeta\prec_{A^{*}}\eta\rangle,$
(91) $\langle\delta(x),\zeta\otimes\eta\rangle=\langle x,\zeta\ast_{A^{*}}\eta\rangle,~\forall~x\in A,\zeta,\eta\in A^{*}.$
Thus, a noncommutative pre-Poisson bialgebra $(A,\succ,\prec,\ast,\Delta_{\succ},\Delta_{\prec},\delta)$ is sometimes denoted by $(A,\succ,\prec,\ast,A^{*},\succ_{A^{*}},\prec_{A^{*}},\ast_{A^{*}})$, where the coherent noncommutative pre-Poisson algebra structure $(\succ_{A^{*}},\prec_{A^{*}},\ast_{A^{*}})$ on the dual space $A^{*}$ corresponds to the coherent noncommutative pre-Poisson coalgebra $(A,\Delta_{\succ},\Delta_{\prec},\delta)$ through Eqs. (90)-(91).
Theorem 3.13.
Let $(A,\succ_{1},\prec_{1},\ast_{1})$ be a coherent noncommutative pre-Poisson algebra equipped with comultiplications $\Delta_{\succ},\Delta_{\prec},\delta:A\longrightarrow A\otimes A$. Suppose that $\Delta_{\succ}^{*},\Delta_{\prec}^{*},\delta^{*}:A^{*}\otimes A^{*}\longrightarrow A^{*}$ induce a coherent noncommutative pre-Poisson algebra structure on $A^{*}$. Put $\succ_{2}=\Delta_{\succ}^{*},\prec_{2}=\Delta_{\prec}^{*},\ast_{2}=\delta^{*}$. Then the following conditions are equivalent:
  1. (a)
    $(A\oplus A^{*},\cdot,[\ ,\ ],\omega)$ is a phase space of the noncommutative Poisson algebra $(A,\cdot_{1},[\ ,\ ]_{1})$.
  2. (b)
    $(A\oplus A^{*},\omega,A,A^{*})$ is a Manin triple of noncommutative pre-Poisson algebras, where the skew-symmetric bilinear form $\omega$ on $A\oplus A^{*}$ is given by Eq. (66).
  3. (c)
    $(A,A^{*},R_{\cdot_{1}}^{*},-L_{\prec_{1}}^{*},-R_{\succ_{1}}^{*},L_{\cdot_{1}}^{*},-\mathrm{ad_{1}}^{*},R^{*}_{\ast_{1}},R_{\cdot_{2}}^{*},-L_{\prec_{2}}^{*},-R_{\succ_{2}}^{*},L_{\cdot_{2}}^{*},-\mathrm{ad_{2}}^{*},R^{*}_{\ast_{2}})$ is a matched pair of coherent noncommutative pre-Poisson algebras.
  4. (d)
    $(A,A^{*},R_{\prec_{1}}^{*},L_{\succ_{1}}^{*},-L_{\ast_{1}}^{*},R_{\prec_{2}}^{*},L_{\succ_{2}}^{*},-L_{\ast_{2}}^{*})$ is a matched pair of coherent noncommutative Poisson algebras.
  5. (e)
    $(A,\succ_{1},\prec_{1},\ast_{1},\Delta_{\succ},\Delta_{\prec},\delta)$ is a noncommutative pre-Poisson bialgebra.
Proof.
[…]
$l_{1}=R_{\prec_{1}}^{*},\ \ \ r_{1}=L_{\succ_{1}}^{*},\ \ \ \rho_{1}=-L_{\ast_{1}}^{*},\ \ \ l_{2}=R_{\prec_{2}}^{*},\ \ \ r_{2}=L_{\succ_{2}}^{*},\ \ \ \rho_{2}=-L_{\ast_{2}}^{*}.$
In fact, we obtain for all $x,y\in A,a,b\in A^{*}$,
$\langle-R_{\prec_{1}}^{*}(L_{\ast_{2}}^{*}(a)x)b,y\rangle$ $=-\langle b,y\prec_{1}(L_{\ast_{2}}^{*}(a)x)\rangle=-\langle L_{\prec_{1}}^{*}(y)b,L_{\ast_{2}}^{*}(a)x\rangle$
$=-\langle a\ast_{2}(L_{\prec_{1}}^{*}(y)b),x\rangle=-\langle(I\otimes L_{\prec_{1}}(y))\delta(x),a\otimes b\rangle,$
$\langle-L_{\succ_{1}}^{*}(L_{\ast_{2}}^{*}(b)x)a,y\rangle$ $=-\langle a,(L_{\ast_{2}}^{*}(b)x)\succ_{1}y\rangle=-\langle R_{\succ_{1}}^{*}(y)a,L_{\ast_{2}}^{*}(b)x\rangle$
$=-\langle b\ast_{2}(R_{\succ_{1}}^{*}(y)a),x\rangle=-\langle(R_{\succ_{1}}(y)\otimes I)\tau\delta(x),a\otimes b\rangle,$
$\langle-(L_{\ast_{1}}^{*}(x)a)\cdot_{2}b,y\rangle=-\langle(L_{\ast_{1}}(x)\otimes I)\Delta(y),a\otimes b\rangle,$
$\langle-a\cdot_{2}L_{\ast_{1}}^{*}(x)b,y\rangle=-\langle(I\otimes L_{\ast_{1}}(x))\Delta(y),a\otimes b\rangle,$
$\langle-L_{\ast_{1}}^{*}(x)(a\cdot_{2}b),y\rangle=-\langle a\cdot_{2}b,x\ast_{1}y\rangle=-\langle\delta(x\ast_{1}y),a\otimes b\rangle.$
[…]
Let $(A,\succ_{A},\prec_{A},\ast_{A},\Delta_{\succ},\Delta_{\prec},\delta)$ be a noncommutative pre-Poisson bialgebra. By Proposition 2.20 and Theorem 3.13, we know that $(D=A\oplus A^{*},\succ_{D},\prec_{D},\ast_{D})$ is a coherent noncommutative pre-Poisson algebra, where
(92) $(x+a)\succ_{D}(y+b)=$ $x\succ_{A}y+(R_{\succ_{A^{*}}}^{*}+R_{\prec_{A^{*}}}^{*})(a)y-L_{\prec_{A^{*}}}^{*}(b)x$
$+a\succ_{A^{*}}b+(R_{\succ_{A}}^{*}+R_{\prec_{A}}^{*})(x)b-L_{\prec_{A}}^{*}(y)a,$
(93) $(x+a)\prec_{D}(y+b)=$ $x\prec_{A}y-R_{\succ_{A^{*}}}^{*}(a)y+(L_{\succ_{A^{*}}}^{*}+L_{\prec_{A^{*}}}^{*})(b)x$
$+a\prec_{A^{*}}b-R_{\succ_{A}}^{*}(x)b+(L_{\succ_{A}}^{*}+L_{\prec_{A}}^{*})(y)a,$
(94) $(x+a)\ast_{D}(y+b)=$ $x\ast_{A}y+(R^{*}_{\ast_{A^{*}}}-L_{\ast_{A^{*}}}^{*})(a)y+R^{*}_{\ast_{A^{*}}}(b)x$
$+a\ast_{A^{*}}b+(R^{*}_{\ast_{A}}-L_{\ast_{A}}^{*})(x)b+R^{*}_{\ast_{A}}(y)a,$
for all $x,y\in A,a,b\in A^{*}$. $(D=A\oplus A^{*},\succ_{D},\prec_{D},\ast_{D})$ is called the double noncommutative pre-Poisson algebra.
3.3. Coboundary noncommutative pre-Poisson bialgebras and the noncommutative pre-Poisson Yang-Baxter equation
In this section, we consider the coboundary noncommutative pre-Poisson bialgebras.
Definition 3.14.
A noncommutative pre-Poisson bialgebra $(A,\succ,\prec,\ast,\Delta_{\succ,r},\Delta_{\prec,r},\delta_{r})$ is called coboundary if there is some $r\in A\otimes A$ such that Eqs. (67)-(68) and (78) hold.
Proposition 3.15.
Let $(A,\succ,\prec,\ast,\Delta_{\succ,r},\Delta_{\prec,r},\delta_{r})$ be a noncommutative pre-Poisson bialgebra and $r=\sum\limits_{i}a_{i}\otimes b_{i}\in A\otimes A$, where $\Delta_{\succ,r},\Delta_{\prec,r},\delta_{r}:A\longrightarrow A\otimes A$ are given by Eqs. (67)-(68) and (78) respectively. Then
  1. (a)
    Eq. (82) holds if and only if
    (95) $(L_{\ast}(x)\otimes I\otimes I)(r_{32}\prec r_{12}-r_{13}\cdot r_{32}+r_{21}\succ r_{31})$
    $-(I\otimes R_{\prec}(x)\otimes I)(r_{12}\ast r_{13}+[r_{32},r_{13}]-r_{12}\ast r_{32})$
    $+(I\otimes I\otimes L_{\cdot}(x))(r_{31}\ast r_{12}+[r_{32},r_{12}]-r_{32}\ast r_{21}-[r_{31},r_{21}]+[r_{32},r_{13}])$
    $+\sum_{i}(L_{\ast}(b_{i}\prec x)\otimes I\otimes I+I\otimes L_{\ast}(b_{i}\prec x)\otimes I)((\tau(r)-r)\otimes a_{i})$
    $+(L_{\ast}(x\cdot a_{i})\otimes I\otimes I)(\tau\otimes I)(b_{i}\otimes r)-(L_{\ast}(x\cdot b_{i})\otimes I\otimes I)(\tau\otimes I)(a_{i}\otimes\tau(r))=0.$
  2. (b)
    Eq. (83) holds if and only if
    (96) $(R_{\cdot}(x)\otimes I\otimes I)(r_{12}\ast r_{23}+[r_{13},r_{23}]-r_{13}\ast r_{32}+[r_{32},r_{12}]+[r_{13},r_{21}])$
    $+(I\otimes L_{\ast}(x)\otimes I)(r_{32}\succ r_{13}-r_{13}\cdot r_{21}+r_{21}\prec r_{23})$
    $+(I\otimes I\otimes L_{\succ}(x))(r_{23}\ast r_{13}-r_{23}\ast r_{21}-[r_{13},r_{21}])$
    $+\sum_{i}(I\otimes L_{\ast}(x\succ b_{i})\otimes I+I\otimes I\otimes L_{\ast}(x\succ b_{i}))(a_{i}\otimes(\tau(r)-r))$
    $+(I\otimes R_{\ast}(x\succ a_{i})\otimes I)(b_{i}\otimes r)-(I\otimes R_{\ast}(x\succ b_{i})\otimes I)(a_{i}\otimes\tau(r))=0.$
  3. (c)
    Eq. (84) holds if and only if
    (97) $(R_{\prec}(x)\otimes I\otimes I)(r_{21}\ast r_{23}+[r_{31},r_{23}]-r_{12}\ast r_{13})$
    $+(I\otimes L_{\succ}(x)\otimes I)(r_{12}\ast r_{13}+[r_{32},r_{13}]-r_{21}\ast r_{23})$
    $+(I\otimes I\otimes\mathrm{ad}(x))(r_{23}\cdot r_{21}-r_{21}\prec r_{13}-r_{23}\succ r_{12}+r_{12}\cdot r_{13}-r_{13}\cdot r_{32})$
    $+\sum_{i}(R_{\prec}(x\ast a_{i})\otimes I\otimes I)((r-\tau(r))\otimes b_{i})+(I\otimes L_{\succ}(x\ast a_{i})\otimes I)((\tau(r)-r)\otimes b_{i})$
    $+(I\otimes I\otimes\mathrm{ad}(x\cdot b_{i}))(a_{i}\otimes\tau(r))-(I\otimes I\otimes\mathrm{ad}(x\cdot a_{i}))(b_{i}\otimes r)=0.$
  4. (d)
    Eq. (85) holds if and only if
    (98) $(L_{\ast}(x)\otimes I\otimes I)(r_{32}\prec r_{12}-r_{13}\cdot r_{32}+r_{21}\succ r_{31})$
    $+(I\otimes L_{\ast}(x)\otimes I)(r_{21}\succ r_{31}-r_{31}\cdot r_{23}+r_{32}\prec r_{12})$
    $+(I\otimes I\otimes\mathrm{ad}(x))(r_{23}\cdot r_{21}-r_{21}\prec r_{13}-r_{23}\succ r_{12}+r_{12}\cdot r_{13}-r_{13}\cdot r_{32})$
    $+\sum_{i}(R_{\prec}(x\ast a_{i})\otimes I\otimes I)((r-\tau(r)\otimes b_{i})$
    $+(R_{\succ}(x\ast a_{i})\otimes I\otimes I)(r\otimes b_{i})-(R_{\succ}(x\ast b_{i})\otimes I\otimes I)(\tau(r)\otimes a_{i})$
    $+(I\otimes L_{\prec}(x\ast a_{i})\otimes I)(\tau(r)\otimes b_{i})-(I\otimes L_{\prec}(x\ast b_{i})\otimes I)(r\otimes a_{i})$
    $+(I\otimes I\otimes R_{\cdot}([x,a_{i}])(\tau\otimes I)(b_{i}\otimes r)-(I\otimes I\otimes R_{\cdot}([x,b_{i}])(\tau\otimes I)(a_{i}\otimes\tau(r))=0.$
  5. (e)
    Eq. (86) holds if and only if
    (99) $(I\otimes L_{\succ}(x\ast y)-I\otimes L_{\ast}(x)L_{\succ}(y)-I\otimes L_{\prec}(y)\mathrm{ad}(x)+L_{\ast}(x)R_{\cdot}(y)\otimes I$
    $-R_{\cdot}(x\ast y)\otimes I+R_{\cdot}(y)\otimes L_{\ast}(x)-L_{\ast}(x)\otimes L_{\cdot}(y))(r-\tau(r))=0.$
  6. (f)
    Eq. (87) holds if and only if
    (100) $(I\otimes R_{\prec}(x)L_{\ast}(y)-I\otimes L_{\ast}(y\prec x)-\mathrm{ad}(y\prec x)\otimes I-L_{\ast}(x)R_{\prec}(y)\otimes I$
    $+\mathrm{ad}(y)\otimes R_{\prec}(x)+L_{\ast}(x)\otimes L_{\cdot}(y))(r-\tau(r))=0.$
  7. (g)
    Eq. (88) holds if and only if
    (101) $(L_{\ast}(x)\otimes R_{\cdot}(y)+L_{\ast}(y)\otimes L_{\cdot}(y)-L_{\ast}(x\cdot y)\otimes I)(r-\tau(r))=0.$
  8. (h)
    Eq. (89) holds if and only if
    (102) $(R_{\cdot}(y)\otimes L_{\ast}(x)-\mathrm{ad}(x)\otimes L_{\succ}(y)+\mathrm{ad}(x)R_{\cdot}(y)\otimes I-R_{\cdot}([x,y])\otimes I-I\otimes L_{\succ}(y)L_{\ast}(x))(r-\tau(r))=0.$
Proof.
Combining Eqs. (67)-(68) and (78), we have for all $x\in A$,
$(\delta_{r}\otimes I)\Delta_{\succ,r}(x)-(\tau\otimes I)(\delta_{r}\otimes I)\Delta_{\succ,r}(x)-(I\otimes\Delta_{\succ,r})\delta_{r}(x)+(\tau\otimes I)(I\otimes\delta)\Delta_{\succ,r}(x)$
$=$ $\sum_{i,j}b_{i}\ast a_{j}\otimes b_{j}\otimes x\cdot a_{i}+a_{j}\otimes[b_{i},b_{j}]\otimes x\cdot a_{i}-(b_{i}\prec x)\ast a_{j}\otimes b_{j}\otimes a_{i}-a_{j}\otimes[b_{i}\prec x,b_{j}]\otimes a_{i}$
$+b_{j}\otimes(b_{i}\prec x)\ast a_{j}\otimes a_{i}+[b_{i}\prec x,b_{j}]\otimes a_{j}\otimes a_{i}-b_{j}\otimes b_{i}\ast a_{j}\otimes x\cdot a_{i}-[b_{i},b_{j}]\otimes a_{j}\otimes x\cdot a_{i}$
$+x\ast a_{i}\otimes b_{j}\prec b_{i}\otimes a_{j}-x\ast a_{i}\otimes b_{j}\otimes b_{i}\cdot a_{j}+a_{i}\otimes b_{j}\prec[x,b_{i}]\otimes a_{j}-a_{i}\otimes b_{j}\otimes[x,b_{i}]\cdot a_{j}$
$-a_{i}\ast a_{j}\otimes b_{i}\prec x\otimes b_{j}-a_{j}\otimes b_{i}\prec x\otimes[a_{i},b_{j}]+(x\cdot a_{i})\ast a_{j}\otimes b_{i}\otimes b_{j}+a_{j}\otimes b_{i}\otimes[x\cdot a_{i},b_{j}]$
$=$ $A(1)+A(2)+A(3),$
where
$A(1)=$ $\sum_{i,j}-(b_{i}\prec x)\ast a_{j}\otimes b_{j}\otimes a_{i}+[b_{i}\prec x,b_{j}]\otimes a_{j}\otimes a_{i}$
$+x\ast a_{i}\otimes b_{j}\prec b_{i}\otimes a_{j}-x\ast a_{i}\otimes b_{j}\otimes b_{i}\cdot a_{j}+(x\cdot a_{i})\ast a_{j}\otimes b_{i}\otimes b_{j},$
$A(2)=$ $\sum_{i,j}-a_{j}\otimes[b_{i}\prec x,b_{j}]\otimes a_{i}+b_{j}\otimes(b_{i}\prec x)\ast a_{j}\otimes a_{i}$
$+a_{i}\otimes b_{j}\prec[x,b_{i}]\otimes a_{j}-a_{i}\ast a_{j}\otimes b_{i}\prec x\otimes b_{j}-a_{j}\otimes b_{i}\prec x\otimes[a_{i},b_{j}],$
$A(3)=$ $\sum_{i,j}b_{i}\ast a_{j}\otimes b_{j}\otimes x\cdot a_{i}+a_{j}\otimes[b_{i},b_{j}]\otimes x\cdot a_{i}-b_{j}\otimes b_{i}\ast a_{j}\otimes x\cdot a_{i}$
$-[b_{i},b_{j}]\otimes a_{j}\otimes x\cdot a_{i}-a_{i}\otimes b_{j}\otimes[x,b_{i}]\cdot a_{j}+a_{j}\otimes b_{i}\otimes[x\cdot a_{i},b_{j}].$
Using Eqs. (3)-(6) and (1), we have
$A(1)=$ $(L_{\ast}(b_{i}\prec x)\otimes I\otimes I)((\tau(r)-r)\otimes a_{i})-b_{j}\ast(b_{i}\prec x)\otimes a_{j}\otimes a_{i}$
$+(L_{\ast}(x)\otimes I\otimes I)(r_{32}\prec r_{12}-r_{13}\cdot r_{32})+a_{i}\ast(a_{j}\prec x)\otimes b_{i}\otimes b_{j}+x\ast(a_{i}\succ a_{j})\otimes b_{i}\otimes b_{j}$
$=$ $(L_{\ast}(b_{i}\prec x)\otimes I\otimes I)((\tau(r)-r)\otimes a_{i})-(R_{\ast}(b_{i}\prec x)\otimes I\otimes I)(\tau(r)\otimes a_{i})$
$+(L_{\ast}(x)\otimes I\otimes I)(r_{32}\prec r_{12}-r_{13}\cdot r_{32}+r_{12}\succ r_{13})+(R_{\ast}(a_{i}\prec x)\otimes I\otimes I)(r\otimes b_{i})$
$A(2)=$ $\sum_{i,j}a_{j}\otimes b_{j}\ast(b_{i}\prec x)\otimes a_{i}+(I\otimes L_{\ast}(b_{i}\prec x)\otimes I)((\tau(r)-r)\otimes a_{i})+a_{i}\otimes(b_{i}\ast b_{j})\otimes\otimes a_{j}$
$-a_{i}\otimes b_{j}\prec[x,b_{i}]\otimes a_{j}-(I\otimes R_{\prec}(x)\otimes I)(r_{12}\ast r_{13}+[r_{32},r_{13}])$
$=$ $(I\otimes L_{\ast}(b_{i}\prec x)\otimes I)((\tau(r)-r)\otimes a_{i})-(I\otimes R_{\prec}(x)\otimes I)(r_{12}\ast r_{13}+[r_{32},r_{13}]-r_{12}\ast r_{32}),$
$A(3)=(I\otimes I\otimes L_{\cdot}(x))(r_{31}\ast r_{12}+[r_{32},r_{12}]-r_{32}\ast r_{21}-[r_{31},r_{21}]-[r_{13},r_{32}]).$
Therefore, Eq. (82) holds if and only if Eq. (95) holds. The remaining Items can be checked in the same way. ∎
Theorem 3.16.
Let $(A,\succ,\prec,\ast)$ be a coherent noncommutative pre-Poisson algebra equipped with linear maps $\Delta_{\succ,r},\Delta_{\prec,r},\delta_{r}:A\longrightarrow A\otimes A$ given by Eqs. (67)-(68) and (78) respectively. Then $(A,\succ,\prec,\ast,\Delta_{\succ,r},\Delta_{\prec,r},\delta_{r})$ is a noncommutative pre-Poisson bialgebra if and only if Eqs. (69)-(73), (79)-(80) and (95)-(102) hold.
Proof.
This follows from Theorem 3.2, Theorem 3.6 and Proposition 3.15.∎
Denote
$D_{1}(r)=r_{23}\succ r_{13}-r_{13}\cdot r_{12}+r_{12}\prec r_{23},$
$D_{2}(r)=r_{13}\cdot r_{23}-r_{12}\succ r_{13}-r_{23}\prec r_{12},$
$S_{1}(r)=r_{23}\ast r_{12}+[r_{13},r_{12}]-r_{23}\ast r_{13}.$
The following conclusion is apparent.
Proposition 3.17.
Let $(A,\succ,\prec,\ast)$ be a coherent noncommutative pre-Poisson algebra and $r=\sum\limits_{i}a_{i}\otimes b_{i}\in A\otimes A$ symmetric. Define linear maps $\Delta_{\succ,r},\Delta_{\prec,r},\delta_{r}:A\longrightarrow A\otimes A$ by Eqs. (67)-(68) and (78) respectively Then $(A,\succ,\prec,\ast,\Delta_{\succ,r},\Delta_{\prec,r},\delta_{r})$. is a noncommutative pre-Poisson bialgebra if and only if Eqs. (74)-(76), (81) and the following equations hold:
(103) $(L_{\ast}(x)\otimes I\otimes I)D_{2}(r)+(I\otimes R_{\prec}(x)\otimes I-I\otimes I\otimes L_{\cdot}(x))S(r)=0,$
(104) $(R_{\cdot}(x)\otimes I\otimes I-I\otimes I\otimes L_{\succ}(x))S_{1}(r)+(I\otimes L_{\ast}(x)\otimes I)D_{1}(r)=0,$
(105) $(R_{\prec}(x)\otimes I\otimes I-I\otimes L_{\succ}(x)\otimes I)S(r)-(I\otimes I\otimes\mathrm{ad}(x))D_{2}(r)=0,$
(106) $(L_{\ast}(x)\otimes I\otimes I+I\otimes L_{\ast}(x)\otimes I+I\otimes I\otimes\mathrm{ad}(x))D_{2}(r)=0.$
Remark 3.18.
Let $\sigma_{132},\sigma_{23},\sigma_{13}:A\otimes A\otimes A\longrightarrow A\otimes A\otimes A$ be linear maps defined by
$\sigma_{132}(x\otimes y\otimes z)=z\otimes x\otimes y,\ \ \ \sigma_{23}(x\otimes y\otimes z)=x\otimes z\otimes y,\ \ \ \sigma_{13}(x\otimes y\otimes z)=z\otimes y\otimes x$
for all $x,y,z\in A$. If $r$ is symmetric, then
$\sigma_{13}S(r)=S_{1}(r),\ \ \ \sigma_{23}D(r)=-D_{1}(r),\ \ \ \sigma_{132}D(r)=D_{2}(r).$
Theorem 3.19.
Let $(A,\succ,\prec,\ast)$ be a coherent noncommutative pre-Poisson algebra and $r\in A\otimes A$ be symmetric. Then $(A,\succ,\prec,\ast,\Delta_{\succ,r},\Delta_{\prec,r},\delta_{r})$ is a noncommutative pre-Poisson bialgebra if $S(r)=D(r)=0$, where $\Delta_{\succ,r},\Delta_{\prec,r},\delta_{r}$ are defined by Eqs. (67)-(68) and (78) respectively.
Proof.
The statement follows from Theorem 3.4, Theorem 3.8, Theorem 3.16, Proposition 3.17 and Remark 3.18. ∎
Definition 3.20.
Let $(A,\succ,\prec,\ast)$ be a coherent noncommutative pre-Poisson algebra and $r\in A\otimes A$. We say that $r$ satisfies the noncommutative-pre-Poisson Yang-Baxter equation or NPP-YBE in short if $r$ satisfies both the D-equation
$D(r)=r_{12}\cdot r_{13}-r_{23}\succ r_{12}-r_{13}\prec r_{23}=0$
and the S-equation:
$S(r)=r_{12}\ast r_{23}-r_{12}\ast r_{13}+[r_{13},r_{23}]=0,$
where $x\cdot y=x\succ y+x\prec y$ and $x,y]=x\ast y-y\ast x$.
Example 3.21.
[…]
Consequently, we arrive at the following conclusion.
Corollary 3.22.
Let $(A,\succ,\prec,\ast)$ be a coherent noncommutative pre-Poisson algebra and $r\in A\otimes A$ be a symmetric solution of the NPP-YBE in $(A,\succ,\prec,\ast)$. Then $(A,\succ,\prec,\ast,\Delta_{\succ,r},\Delta_{\prec,r},\delta_{r})$ is a noncommutative pre-Poisson bialgebra, where $\Delta_{\succ,r},\Delta_{\prec,r},\delta_{r}:A\rightarrow A\otimes A$ are linear maps given by Eqs. (67)-(68) and (78) respectively.
For a vector space $A$, the isomorphism $A\otimes A^{*}\simeq Hom(A^{*},A)$ identifies an element $r\in A\otimes A$ with a map $T_{r}:A^{*}\longrightarrow A$. Explicitly,
(107) $T_{r}:A^{*}\longrightarrow A,\ \ \ \langle T_{r}(\zeta),\eta\rangle=\langle r,\zeta\otimes\eta\rangle,\ \ \ \forall~\zeta,\eta\in A^{*}.$
It is clearly that $T_{r}^{*}=T_{\tau(r)},~T_{r-\tau(r)}^{*}=-T_{r-\tau(r)}$.
Definition 3.23.
Assume that $(A,\succ,\prec,\ast)$ is a noncommutative pre-Poisson algebra and $(V,$ $\succ_{V},\prec_{V},\ast_{V},l_{\succ},r_{\succ},l_{\prec},r_{\prec},l_{\ast},r_{\ast})$ is an A-noncommutative pre-Poisson algebra. A relative Rota-Baxter operator $T$ of weight $\lambda$ on $(A,\succ,\prec,\ast)$ associated to $(V,\succ_{V},\prec_{V},\ast_{V},l_{\succ},r_{\succ},l_{\prec},r_{\prec},l_{\ast},r_{\ast})$ is a linear map $T:V\longrightarrow A$ satisfying
$T(u)\succ T(v)=T(l_{\succ}(T(u))v+r_{\succ}(T(v))u+\lambda u\succ_{V}v),$
$T(u)\prec T(v)=T(l_{\prec}(T(u))v+r_{\prec}(T(v))u+\lambda u\prec_{V}v),$
$T(u)\ast T(v)=T(l_{\ast}(T(u))v+r_{\ast}(T(v))u+\lambda u\ast_{V}v).$
When $u\succ_{V}v=u\prec_{V}v=u\ast_{V}v=0$ for all $u,v\in V$, then $T$ is simply a relative Rota-Baxter operator ( $\mathcal{O}$-operator) on $(A,\succ,\prec,\ast)$ associated to the representation $(V,l_{\succ},r_{\succ},l_{\prec},r_{\prec},l_{\ast},r_{\ast})$.
Combining Theorem 4.4.11 [4] and Theorem 6.6 [3], we have the following conclusion.
Theorem 3.24.
Let $(A,\succ,\prec,\ast)$ be a coherent noncommutative pre-Poisson algebra and $r\in A\otimes A$ be symmetric. Then the following conditions are equivalent:
  1. (a)
    $r$ is a solution of the NPP-YBE in $(A,\succ,\prec,\ast)$.
  2. (b)
    $T_{r}$ is an $\mathcal{O}$-operator of the noncommutative Poisson algebra $(A,\cdot,[\ ,\ ])$ associated to $(A^{*},R_{\prec}^{*},L_{\succ}^{*},-L_{\ast}^{*})$.
  3. (c)
    $T_{r}$ is an $\mathcal{O}$-operator of $(A,\succ,\prec,\ast)$ associated to $(A^{*},R_{\cdot}^{*},-L_{\prec}^{*},-R_{\succ}^{*},L_{\cdot}^{*},-\mathrm{ad}^{*},R^{*}_{\ast})$.
Theorem 3.25.
Let $(V,l_{\succ},r_{\succ},l_{\prec},r_{\prec},l_{\ast},r_{\ast})$ be a representation of a coherent noncommutative pre-Poisson algebra satisfying Eqs. (18)-(20). Assume that $(V^{*},r_{\cdot}^{*},-l_{\prec}^{*},-r_{\succ}^{*},l_{\cdot}^{*},r_{\ast}^{*}-l_{\ast}^{*},r_{\ast}^{*})$ is the dual representation of $A$ given by Proposition 2.6. Denote $\hat{A}=A\ltimes V^{*}$ and suppose that $T:V\longrightarrow A$ is a linear map which is identifies an element in $\hat{A}\otimes\hat{A}$ through ($Hom(V,A)\simeq A\otimes V^{*}\subseteq\hat{A}\otimes\hat{A}$). Then $r=T+\tau(T)$ is a symmetric solution of the NPP-YBE in the coherent noncommutative pre-Poisson algebra $\hat{A}$ if and only if $T$ is an $\mathcal{O}$-operator on $(A,\succ,\prec,\ast)$ associated to $(V,l_{\succ},r_{\succ},l_{\prec},r_{\prec},l_{\ast},r_{\ast})$.
Proof.
Based on Remark 2.11, if $(V,l_{\succ},r_{\succ},l_{\prec},r_{\prec},l_{\ast},r_{\ast})$ is a representation of a coherent noncommutative pre-Poisson algebra $(A,\succ,\prec,\ast)$ satisfying Eqs. (18)-(20), then $A\ltimes V^{*}$ is a coherent noncommutative pre-Poisson algebra. Combining Theorem 4.4.13 [4] and Theorem 6.9 [3], we obtain the desired result. ∎
Example 3.26.
Let $(A=ke_{1}\oplus ke_{2},\succ,\prec,\ast)$ be the 2-dimensional coherent noncommutative pre-Poisson algebra given in Example 2.3. Define a linear map $T:A\longrightarrow A$ by $T(e_{1})=e_{2},~T(e_{2})=0$. It is easy to check that $T$ an $\mathcal{O}$-operator of $(A,\succ,\prec,\ast)$ associated to the regular representation $(A,L_{\succ},R_{\succ},L_{\prec},R_{\prec},L_{\ast},R_{\ast})$. Denote the dual basis of $A^{*}$ by $\{e_{1}^{*},e_{2}^{*}\}$. By Remark 2.11, the semi-direct product $A\ltimes A^{*}$ with respect to its representation $(A^{*},R_{\cdot}^{*},-L_{\prec}^{*},-R_{\succ}^{*},L_{\cdot}^{*},-\mathrm{ad}^{*},R^{*}_{\ast})$ is a coherent noncommutative pre-Poisson algebra with the multiplications $(\succ,\prec,\ast)$ defined by
$(x+u)\succ(y+v)=x\succ y+R_{\cdot}^{*}(x)v-L_{\prec}^{*}(y)u,$
$(x+u)\prec(y+v)=x\prec y-R_{\succ}^{*}(x)v+L_{\cdot}^{*}(y)u,$
$(x+u)\ast(y+v)=x\ast y-\mathrm{ad}^{*}(x)v+R_{\ast}^{*}(y)u,\ \ ~\forall~x,y\in A,u,v\in A^{*}.$
Explicitly, the non-trivial multiplications $(\succ,\prec,\ast)$ are given by
$e_{1}\succ e_{2}=-e_{1}\prec e_{2}=e_{2}\prec e_{1}=e_{2}\ast e_{1}=e_{2},\ \ \ \ \ e_{1}\succ e_{1}=e_{1}\ast e_{1}=e_{1},$
$e_{2}^{*}\succ e_{2}=-e_{1}\succ e_{1}^{*}=e_{2}^{*}\prec e_{2}=e_{1}\prec e_{1}^{*}=e_{1}^{*}\prec e_{1}=e_{2}\ast e_{2}^{*}=-e_{1}^{*}\ast e_{1}=e_{1}^{*}$
$e_{1}\succ e_{2}^{*}=e_{2}^{*}\succ e_{1}=e_{1}\ast e_{2}^{*}=e_{2}^{*}\ast e_{1}=-e_{2}^{*},$
In view of Theorem 3.25, we know that $r=\sum_{i=1}^{2}T(e_{i})\otimes e_{i}^{*}+e_{i}^{*}\otimes T(e_{i})$ is a symmetric solution of the NPP-YBE in the coherent noncommutative pre-Poisson algebra $A\ltimes A^{*}$. According to Corollary 3.22, $(A\ltimes A^{*},\Delta_{\succ},\Delta_{\prec},\delta)$ is a noncommutative pre-Poisson bialgebra with the linear maps $\Delta_{\succ},\Delta_{\prec},\delta:A\ltimes A^{*}\longrightarrow(A\ltimes A^{*})\otimes(A\ltimes A^{*})$ defined respectively by
$\Delta_{\succ}(x)=(I\otimes L_{\cdot}(x)-R_{\prec}(x)\otimes I)\tau(r),$
$\Delta_{\prec}(x)=(R_{\cdot}(x)\otimes I-I\otimes L_{\succ}(x))r,$
$\delta(x)=-(L_{\ast}(x)\otimes I+I\otimes\mathrm{ad}(x))r,~~\forall~x\in A\ltimes A^{*}.$
Explicitly, the comultiplications $(\Delta_{\succ},\Delta_{\prec},\delta)$ are given as follows (only non-zero operations are listed):
$\delta(e_{1})=e_{1}^{*}\otimes e_{2}-e_{2}\otimes e_{1}^{*},\ \ \ \delta(e_{2}^{*})=e_{1}^{*}\otimes e_{1}^{*},$
$\Delta_{\succ}(e_{1})=-e_{2}\otimes e_{1}^{*}-e_{1}^{*}\otimes e_{2},\ \ \ \Delta_{\succ}(e_{2}^{*})=e_{1}^{*}\otimes e_{1}^{*},$
$\Delta_{\prec}(e_{1})=e_{2}\otimes e_{1}^{*}-2e_{1}^{*}\otimes e_{2}.$
4. Quasi-triangular noncommutative pre-Poisson bialgebras and factorizable noncommutative pre-Poisson bialgebras
Having shown in Theorem 3.19 that symmetric solutions of the NPP-YBE yield coboundary noncommutative pre-Poisson bialgebras, we now address the more general case of non-symmetric solutions.
4.1. Quasi-triangular noncommutative pre-Poisson bialgebras
Definition 4.1.
Let $(A,\succ,\prec,\ast)$ be a noncommutative pre-Poisson algebra and $r\in A\otimes A$. Then $r$ is called invariant if
(108) $(I\otimes L_{\cdot}(x)-R_{\prec}(x)\otimes I)\tau(r)=0,$
(109) $(R_{\cdot}(x)\otimes I-I\otimes L_{\succ}(x))r=0,$
(110) $(L_{\ast}(x)\otimes I+I\otimes\mathrm{ad}(x))r=0,~\forall~x\in A.$
Lemma 4.2.
Let $(A,\succ,\prec,\ast)$ be a noncommutative pre-Poisson algebra and $r\in A\otimes A$. Then $r$ is invariant if and only if
(111) $R_{\prec}(x)T_{r}(\zeta)=T_{r}(L_{\cdot}^{*}(x)\zeta),$
(112) $L_{\succ}(x)T_{r}(\zeta)=T_{r}(R_{\cdot}^{*}(x)\zeta),$
(113) $\mathrm{ad}(x)T_{r}(\zeta)=-T_{r}(L_{\ast}^{*}(x)\zeta),~~\forall~x\in A,\zeta\in A^{*}.$
Moreover, Eqs. (108)-(110) hold if and only if the following equations hold:
(114) $R_{\cdot}^{*}(T_{\tau(r)}(\zeta))\eta=L_{\prec}^{*}(T_{r}(\eta))\zeta,$
(115) $L_{\cdot}^{*}(T_{\tau(r)}(\eta))\zeta=R_{\succ}^{*}(T_{r}(\zeta))\eta$
(116) $\mathrm{ad}^{*}(T_{r}(\zeta))\eta=R_{\ast}^{*}(T_{\tau(r)}(\eta))\zeta,~~\forall~x\in A,\zeta\in A^{*}.$
Proof.
Based on Lemma 2.12 [30] and Lemma 2.9 [29], Eqs. (111)-(113) hold. For all $x\in A,\zeta,\eta\in A^{*}$, we have
$\langle(I\otimes L_{\cdot}(x)-R_{\prec}(x)\otimes I)\tau(r),\zeta\otimes\eta\rangle=$ $\langle\tau(r),\zeta\otimes L_{\cdot}^{*}(x)\eta\rangle-\langle\tau(r),R_{\prec}^{*}(x)\zeta\otimes\eta\rangle$
$=$ $\langle T_{\tau(r)}(\zeta),L_{\cdot}^{*}(x)\eta\rangle-\langle T_{r}(\eta),R_{\prec}^{*}(x)\zeta\rangle$
$=$ $\langle x\cdot T_{\tau(r)}(\zeta),\eta\rangle-\langle T_{r}(\eta)\prec x,\zeta\rangle$
$=$ $\langle R_{\cdot}^{*}(T_{\tau(r)}(\zeta))\eta-L_{\prec}^{*}(T_{r}(\eta))\zeta,x\rangle.$
[…]
Proposition 4.3.
Let $(A,\succ,\prec,\ast)$ be a noncommutative pre-Poisson algebra and $r\in A\otimes A$. Then the following conditions are equivalent:
  1. (a)
    $r-\tau(r)$ is invariant.
  2. (b)
    The following equations hold:
    (117) $R_{\cdot}^{*}(T_{r-\tau(r)}(\zeta))\eta=-L_{\prec}^{*}(T_{r-\tau(r)}(\eta))\zeta,\ \ L_{\cdot}^{*}(T_{r-\tau(r)}(\eta))\zeta=-R_{\succ}^{*}(T_{r-\tau(r)}(\zeta))\eta,\ \ \mathrm{ad}^{*}(T_{r-\tau(r)}(\zeta))\eta=-R_{\ast}^{*}(T_{r-\tau(r)}(\eta))\zeta.$
  3. (c)
    The following equations hold:
    (118) $R_{\prec}(x)T_{r-\tau(r)}(\zeta)=T_{r-\tau(r)}(L_{\cdot}^{*}(x)\zeta),\ \ L_{\succ}(x)T_{r-\tau(r)}(\zeta)=T_{r-\tau(r)}(R_{\cdot}^{*}(x)\zeta),\ \ \mathrm{ad}(x)T_{r-\tau(r)}(\zeta)=-T_{r-\tau(r)}(L_{\ast}^{*}(x)\zeta).$
  4. (d)
    The following equations hold:
    (119) $T_{r-\tau(r)}(R_{\prec}^{*}x)(\zeta)=L_{\cdot}(x)T_{r-\tau(r)}(\zeta),\ \ T_{r-\tau(r)}(L_{\succ}^{*}(x)\zeta)=R_{\cdot}(x)T_{r-\tau(r)}(\zeta),\ \ T_{r-\tau(r)}(\mathrm{ad}^{*}(x)\zeta)=-L_{\ast}(x)T_{r-\tau(r)}(\zeta).$
for all $x\in A,~\zeta,\eta\in A^{*}$
Proof.
It follows from Lemma 2.13 [30] and Proposition 2.12 [29] and Lemma 4.2. ∎
By Eqs. (117)-(119), we have
(120) $T_{r-\tau(r)}(R_{\succ}^{*}(x)\zeta)=-L_{\prec}(x)T_{r-\tau(r)}(\zeta),\ \ \ T_{r-\tau(r)}(L_{\prec}^{*}(x)\zeta)=-R_{\succ}(x)T_{r-\tau(r)}(\zeta),$
(121) $T_{r-\tau(r)}(R_{\ast}^{*}(x)\zeta)=R_{\ast}(x)T_{r+\tau(r)}(\zeta),$
(122) $L_{\ast}^{*}(T_{r-\tau(r)}(\zeta))\eta=-L_{\ast}^{*}(T_{r-\tau(r)}(\eta))\zeta,\ \ \ R_{\prec}^{*}(T_{r-\tau(r)}(\zeta))\eta=L_{\succ}^{*}(T_{r-\tau(r)}(\eta))\zeta.$
Theorem 4.4.
Let $(A,\succ,\prec,\ast)$ be a coherent noncommutative pre-Poisson algebra and $r=\sum_{i}a_{i}\otimes b_{i}\in A\otimes A$. Assume that $\Delta_{\succ,r},\Delta_{\prec,r},\delta_{r}$ are given by Eqs. (67)-(68) and (78). If $r$ is a solution of the NPP-YBE in $(A,\succ,\prec,\ast)$ and $r-\tau(r)$ is invariant. Then $(A,\succ,\prec,\ast,\Delta_{\succ,r},\Delta_{\prec,r},\delta_{r})$ is a noncommutative pre-Poisson bialgebra.
Proof.
Since $r$ is a solution of the NPP-YBE in $(A,\succ,\prec,\ast)$ and $r-\tau(r)$ is invariant, $S(r)=D(r)=0$ and Eqs. (117)-(122) hold. By Proposition 2.16 [30], By Proposition 2.10 [29] and Theorem 2.14 [29], $(A,\succ,\prec,\Delta_{\succ,r},\Delta_{\prec,r})$ is a dendriform bialgebra and $(A,\ast,\delta_{r})$ is a pre-Lie bialgebra. Note that
$r_{32}\prec r_{12}-r_{13}\cdot r_{32}+r_{21}\succ r_{31}$
$=$ $-\sigma_{132}D(r)+(r_{31}-r_{13})\cdot r_{32}+(r_{21}-r_{12})\succ r_{31}$
$=$ $-\sigma_{132}D(r)+\sum_{i}(I\otimes I\otimes R_{\cdot}(a_{i}))(\tau\otimes I)(b_{i}\otimes(\tau(r)-r))+(R_{\succ}(b_{i})\otimes I\otimes I)((\tau(r)-r)\otimes a_{i}),$
$r_{12}\ast r_{13}+[r_{32},r_{13}]-r_{12}\ast r_{32}$
$=$ $-S(r)+r_{12}\ast(r_{23}-r_{32})+[r_{13},r_{23}-r_{32}]$
$=$ $-S(r)+\sum_{i}(I\otimes L_{\ast}(b_{i})\otimes I)(a_{i}\otimes(r-\tau(r)))+(I\otimes I\otimes\mathrm{ad}(b_{i}))(a_{i}\otimes(r-\tau(r))),$
$r_{31}\ast r_{12}+[r_{32},r_{12}]-r_{32}\ast r_{21}-[r_{31},r_{21}]+[r_{32},r_{13}]$
$=$ $-(\sigma_{23}+\sigma_{13})S(r)+r_{32}\ast(r_{13}-r_{31})$
$=$ $-(\sigma_{23}+\sigma_{13})S(r)+\sum_{i}(I\otimes I\otimes L_{\ast}(a_{i})-R_{\ast}(a_{i})\otimes I\otimes I)(\tau\otimes I)(b_{i}\otimes(r-\tau(r))).$
Using Eqs. (118)-(119) and (107), we have for all $\zeta,\eta,\theta\in A^{*}$,
$\langle(I\otimes R_{\prec}(x)L_{\ast}(b_{i})\otimes I+I\otimes R_{\prec}(x)\otimes\mathrm{ad}(b_{i}))(a_{i}\otimes(r-\tau(r))),\zeta\otimes\eta\otimes\theta\rangle$
$=$ $\langle r-\tau(r),L_{\ast}^{*}(b_{i})R_{\prec}^{*}(x)\eta\otimes\theta+R_{\prec}^{*}(x)\eta\otimes\mathrm{ad}^{*}(b_{i})\theta\rangle\langle a_{i},\zeta\rangle$
$=$ $\langle T_{r-\tau(r)}(L_{\ast}^{*}(b_{i})R_{\prec}^{*}(x)\eta)+\mathrm{ad}(b_{i})T_{r-\tau(r)}(R_{\prec}^{*}(x)\eta),\theta\rangle\langle a_{i},\zeta\rangle$
$=$ $\langle-\mathrm{ad}(T_{r}(\zeta))T_{r-\tau(r)}(R_{\prec}^{*}(x)\eta)+\mathrm{ad}(T_{r}(\zeta))T_{r-\tau(r)}(R_{\prec}^{*}(x)\eta),\theta\rangle$
$=$ $0.$
By Eqs. (118), (121), (16) and (107), we get for all $\zeta,\eta,\theta\in A^{*}$,
$\langle(L_{\ast}(x)\otimes I\otimes R_{\cdot}(a_{i})+I\otimes I\otimes L_{\cdot}(x)L_{\ast}(a_{i})-R_{\ast}(a_{i})\otimes I\otimes L_{\cdot}(x))(\tau\otimes I)(b_{i}\otimes(\tau(r)-r)),\zeta\otimes\eta\otimes\theta\rangle$
$=$ $\langle\tau(r)-r,L_{\ast}^{*}(x)\zeta\otimes R_{\cdot}^{*}(a_{i})\theta+\zeta\otimes L_{\ast}^{*}(a_{i})L_{\cdot}^{*}(x)\theta-R_{\ast}^{*}(a_{i})\zeta\otimes L_{\cdot}^{*}(x)\theta\rangle\langle b_{i},\eta\rangle$
$=$ $-\langle R_{\cdot}(a_{i})T_{r-\tau(r)}(L_{\ast}^{*}(x)\zeta)+L_{\cdot}(x)L_{\ast}(a_{i})T_{r-\tau(r)}(\zeta)-L_{\cdot}(x)T_{r-\tau(r)}(R_{\ast}^{*}(a_{i})\zeta),\theta\rangle\langle b_{i},\eta\rangle$
$=$ $\langle R_{\cdot}(a_{i})\mathrm{ad}(x)T_{r-\tau(r)}(\zeta)+L_{\cdot}(x)L_{\ast}(a_{i})T_{r-\tau(r)}(\zeta)-L_{\cdot}(x)R_{\ast}(a_{i})T_{r-\tau(r)}(\zeta),\theta\rangle\langle b_{i},\eta\rangle$
$=$ $\langle R_{\cdot}(a_{i})\mathrm{ad}(x)T_{r-\tau(r)}(\zeta)+L_{\cdot}(x)\mathrm{ad}(a_{i})T_{r-\tau(r)}(\zeta),\theta\rangle\langle b_{i},\eta\rangle$
$=$ $\langle\mathrm{ad}(x\cdot T_{\tau(r)}(\eta))T_{r-\tau(r)}(\zeta),\theta\rangle.$
According to Eqs. (118), (122), (16) and (107), we get for all $\zeta,\eta,\theta\in A^{*}$,
$\langle(L_{\ast}(x\cdot a_{i})\otimes I\otimes I)(\tau\otimes I)(b_{i}\otimes r)-(L_{\ast}(x\cdot b_{i})\otimes I\otimes I)(\tau\otimes I)(a_{i}\otimes\tau(r)),\zeta\otimes\eta\otimes\theta\rangle$
$=$ $\langle b_{i},\eta\rangle\langle L_{\ast}^{*}(x\cdot a_{i})\zeta\otimes\theta,r\rangle-\langle a_{i},\eta\rangle\langle L_{\ast}^{*}(x\cdot b_{i})\zeta\otimes\theta,\tau(r)\rangle$
$=$ $\langle T_{r}(L_{\ast}^{*}(x\cdot T_{\tau(r)}(\eta))\zeta),\theta\rangle-\langle T_{\tau(r)}(L_{\ast}^{*}(x\cdot T_{r}(\eta))\zeta),\theta\rangle$
$=$ $\langle T_{r-\tau(r)}(L_{\ast}^{*}(x\cdot T_{\tau(r)}(\eta))\zeta),\theta\rangle-\langle T_{\tau(r)}(L_{\ast}^{*}(x\cdot T_{r-\tau(r)}(\eta))\zeta),\theta\rangle$
$=$ $-\langle\mathrm{ad}(x\cdot T_{\tau(r)}(\eta)T_{r-\tau(r)}(\zeta),\theta\rangle-\langle T_{\tau(r)}(L_{\ast}^{*}(T_{r-\tau(r)}(R_{\prec}^{*}(x)\eta))\zeta),\theta\rangle$
$=$ $-\langle\mathrm{ad}(x\cdot T_{\tau(r)}(\eta)T_{r-\tau(r)}(\zeta),\theta\rangle+\langle L_{\ast}^{*}(T_{r-\tau(r)}(\zeta))R_{\prec}^{*}(x)\eta,T_{r}(\theta)\rangle$
$=$ $-\langle\mathrm{ad}(x\cdot T_{\tau(r)}(\eta)T_{r-\tau(r)}(\zeta),\theta\rangle+\langle\eta,R_{\prec}(x)L_{\ast}(T_{r-\tau(r)}(\zeta))T_{r}(\theta)\rangle.$
In view of Eqs. (118), (120), (14) and (107), we obatin for all $\zeta,\eta,\theta\in A^{*}$,
$\langle(L_{\ast}(b_{i}\prec x)\otimes I\otimes I+I\otimes L_{\ast}(b_{i}\prec x)\otimes I+L_{\ast}(x)R_{\succ}(b_{i})\otimes I\otimes I)((\tau(r)-r)\otimes a_{i}),\zeta\otimes\eta\otimes\theta\rangle$
$=$ $\langle\tau(r)-r,L_{\ast}^{*}(b_{i}\prec x)\zeta\otimes\eta+\zeta\otimes L_{\ast}^{*}(b_{i}\prec x)\eta+R_{\succ}^{*}(b_{i})L_{\ast}^{*}(x)\zeta\otimes\eta\rangle\langle a_{i},\theta\rangle$
$=$ $-\langle T_{r-\tau(r)}(L_{\ast}^{*}(b_{i}\prec x)\zeta)+L_{\ast}(b_{i}\prec x)T_{r-\tau(r)}(\zeta)+T_{r-\tau(r)}(R_{\succ}^{*}(b_{i})L_{\ast}^{*}(x)\zeta),\eta\rangle\langle a_{i},\theta\rangle$
$=$ $\langle(\mathrm{ad}(b_{i}\prec x)-L_{\ast}(b_{i}\prec x)-L_{\prec}(b_{i})\mathrm{ad}(x))T_{r-\tau(r)}(\zeta),\eta\rangle\langle a_{i},\theta\rangle$
$=$ $\langle-R_{\prec}(x)R_{\ast}(T_{r}(\theta))T_{r-\tau(r)}(\zeta),\eta\rangle.$
Combining the above equalities, we get that Eq. (95) holds. By the same token, Eqs. (96)-(102) hold. Combining Theorem 3.16, we complete the proof. ∎
Example 4.5.
Let $(A=ke_{1}\oplus ke_{2},\succ,\prec,\ast)$ be the 2-dimensional coherent noncommutative pre-Poisson algebra given in Example 2.3. It is easy to check that $r=e_{1}\otimes e_{2}$ is a non-symmetric solution of the NPP-YBE in $(A,\succ,\prec,\ast)$ and $r-\tau(r)=e_{1}\otimes e_{2}-e_{2}\otimes e_{1}$ is invariant.
Let $(A,\succ,\prec,\ast,\Delta_{\succ,r},\Delta_{\prec,r},\delta_{r})$ be a noncommutative pre-Poisson bialgebra and $r\in A\otimes A$, where the comultiplications $\Delta_{\succ,r},\Delta_{\prec,r},\delta_{r}$ are defined by Eqs. (67)-(68) and (78). According to Remark 3.12, there is a coherent noncommutative pre-Poisson algebra structure $(\succ_{r},\prec_{r},\ast_{r})$ on $A^{*}$. By direct computations,
(123) $\zeta\succ_{r}\eta=R_{\cdot}^{*}(T_{\tau(r)}(\zeta))\eta-L_{\prec}^{*}(T_{r}(\eta))\zeta,$
(124) $\zeta\prec_{r}\eta=-R_{\succ}^{*}(T_{r}(\zeta))\eta+L_{\cdot}^{*}(T_{\tau(r)}(\eta))\zeta,$
(125) $\zeta\ast_{r}\eta=\mathrm{ad}^{*}(T_{r}(\zeta))\eta-R_{\ast}^{*}(T_{\tau(r)}(\eta))\zeta,$
for all $\zeta,\eta\in A^{*}$. And the sub-adjacent noncommutative Poisson algebra structure $(\cdot_{r},[\ ,\ ]_{r})$ on $A^{*}$ is given by
$\zeta\cdot_{r}\eta=\zeta\succ_{r}\eta+\zeta\prec_{r}\eta,\ \ \ [\zeta,\eta]_{r}=\zeta\ast_{r}\eta-\eta\ast_{r}\zeta,$
where $\mathrm{ad}=L_{\ast}-R_{\ast},~\cdot=\succ+\prec$.
Definition 4.6.
Let $(A,\succ,\prec,\ast)$ be a coherent noncommutative pre-Poisson algebra and $r\in A\otimes A$. If $r$ is a solution of the NPP-YBE in $(A,\succ,\prec,\ast)$ and $r-\tau(r)$ is invariant, then the noncommutative pre-Poisson bialgebra $(A,\succ,\prec,\ast,\Delta_{\succ,r},\Delta_{\prec,r},\delta_{r})$ induced by $r$ is called a quasi-triangular noncommutative pre-Poisson bialgebra. In particular, $(A,\succ,\prec,\ast,\Delta_{\succ,r},\Delta_{\prec,r},\delta_{r})$ is called a triangular noncommutative pre-Poisson bialgebra if $r$ is symmetric, where $\Delta_{\succ,r},\Delta_{\prec,r}$ and $\delta_{r}$ are given by Eqs. (67)-(68) and (78) respectively.
Combining Theorem 2.18 [30] and Theorem 2.14 [29], we get the following results:
Theorem 4.7.
Let $(A,\succ,\prec,\ast)$ be a coherent noncommutative pre-Poisson algebra and $r\in A\otimes A$. Suppose that $r-\tau(r)$ is invariant. Then the following conditions are equivalent:
  1. (a)
    $r$ is a solution of the NPP-YBE in $(A,\succ,\prec,\ast)$.
  2. (b)
    $(A^{*},\cdot_{r},[\ ,\ ]_{r})$ is a coherent Poisson algebra and the linear maps $T_{r},T_{\tau(r)}$ are both Poisson algebra homomorphisms from $(A^{*},\cdot_{r},[\ ,\ ]_{r})$ to $(A,\cdot,[\ ,\ ])$.
  3. (c)
    $(A^{*},\succ_{r},\prec_{r},\ast_{r})$ is a coherent noncommutative pre-Poisson algebra and the linear maps $T_{r},T_{\tau(r)}$ are noncommutative pre-Poisson algebra homomorphisms from $(A^{*},\succ_{r},\prec_{r},\ast_{r})$ to $(A,\succ,\prec,\ast)$.
Corollary 4.8.
Let $(A,\succ,\prec,\ast,\Delta_{\succ,r},\Delta_{\prec,r},\delta_{r})$ be a quasi-triangular noncommutative pre-Poisson bialgebra. Then
  1. (a)
    $T_{r},T_{\tau(r)}$ are both Poisson algebra homomorphisms from $(A^{*},\cdot_{r},[\ ,\ ]_{r})$ to $(A,\cdot,[\ ,\ ])$.
  2. (b)
    $T_{r},T_{\tau(r)}$ are noncommutative pre-Poisson algebra homomorphisms from $(A^{*},\succ_{r},\prec_{r},\ast_{r})$ to $(A,\succ,\prec,\ast)$.
4.2. Factorizable noncommutative pre-Poisson bialgebras
We begin by introducing the notion of a factorizable noncommutative pre-Poisson bialgebra, which constitutes a key subclass of the quasi-triangular family. We then prove that the double of any noncommutative pre-Poisson bialgebra naturally carries the structure of such a factorizable bialgebra.
Definition 4.9.
A quasi-triangular noncommutative pre-Poisson bialgebra $(A,\succ,\prec,\ast,\Delta_{\succ,r},\Delta_{\prec,r},\delta_{r})$ is called factorizable if the linear map $T_{r-\tau(r)}:A^{*}\longrightarrow A$ given by Eq. (107) is a linear isomorphism of vector spaces, where $\Delta_{\succ,r},\Delta_{\prec,r},\delta_{r}$ are defined by Eqs. (67)-(68) and (78) respectively.
Proposition 4.10.
Let $(A,\succ,\prec,\ast,\Delta_{\succ,r},\Delta_{\prec,r},\delta_{r})$ be a quasi-triangular (factorizable) noncommutative pre-Poisson bialgebra. Then $(A,\succ,\prec,\ast,\Delta_{\succ,\tau(r)},\Delta_{\prec,\tau(r)},\delta_{\tau(r)})$ is also a quasi-triangular (factorizable) noncommutative pre-Poisson bialgebra.
Proof.
Since $(A,\succ,\prec,\ast,\Delta_{\succ,r},\Delta_{\prec,r},\delta_{r})$ is a quasi-triangular (factorizable) noncommutative pre-Poisson bialgebra, $D(r)=S(r)=0$ and $r-\tau(r)$ is invariant. Note that
$D(\tau(r))=r_{21}\cdot r_{31}-r_{32}\succ r_{21}-r_{31}\prec r_{32}$
$=$ $D(r)+(r_{21}-r_{12})\cdot r_{31}+r_{12}\cdot(r_{31}-r_{13})-r_{32}\succ(r_{21}-r_{12})$
$+(r_{23}-r_{32})\succ r_{12}+(r_{13}-r_{31})\prec r_{32}+r_{13}\prec(r_{23}-r_{32})$
$=$ $D(r)-\sum_{i}(R_{\cdot}(b_{i})\otimes I-I\otimes L_{\succ}(b_{i}))(r-\tau(r))\otimes a_{i}+a_{i}\otimes(R_{\succ}(b_{i})\otimes I+I\otimes L_{\prec}(b_{i}))(r-\tau(r))$
$-(L_{\cdot}(a_{i})\otimes I\otimes I-I\otimes I\otimes R_{\prec}(a_{i}))(\tau\otimes I)(b_{i}\otimes(r-\tau(r))$
$=$ $0.$
An analogous argument shows that $S(\tau(r))=0$. The proof is completed. ∎
Proposition 4.11.
Let $(A,\succ,\prec,\ast,\Delta_{\succ,r},\Delta_{\prec,r},\delta_{r})$ be a factorizable noncommutative pre-Poisson bialgebra. Then $\mathrm{Im}(T_{r}\oplus T_{\tau(r)})$ is a noncommutative pre-Poisson subalgebra of the direct sum noncommutative pre-Poisson algebra $A\oplus A$, which is isomorphic to the noncommutative pre-Poisson algebra $(A^{*},\succ_{r},\prec_{r},\ast_{r})$. Furthermore, any $x\in A$ has a unique decomposition $x=x_{1}-x_{2}$, where $(x_{1},x_{2})\in\mathrm{Im}(T_{r}\oplus T_{\tau(r)})$ and
$T_{r}\oplus T_{\tau(r)}:A^{*}\longrightarrow A\oplus A,~~(T_{r}\oplus T_{\tau(r)})(\zeta)=(T_{r}(\zeta),T_{\tau(r)}(\zeta)),~\forall~\zeta\in A^{*}.$
Proof.
In view of Definition 4.9 and Corollary 4.8, $T_{r},T_{\tau(r)}$ are noncommutative pre-Poisson algebra homomorphisms and $T_{r-\tau(r)}$ is a linear isomorphism. It follows that $T_{r}\oplus T_{\tau(r)}$ is a noncommutative pre-Poisson algebra homomorphism and $\mathrm{Ker}(T_{r}\oplus T_{\tau(r)})=0$. Therefore, $\mathrm{Im}(T_{r}\oplus T_{\tau(r)})$ is isomorphic to $(A^{*},\succ_{r},\prec_{r},\ast_{r})$ as noncommutative pre-Poisson algebras. For all $x\in A$,
$x=T_{r-\tau(r)}T_{r-\tau(r)}^{-1}(x)=T_{r}T_{r-\tau(r)}^{-1}(x)-T_{\tau(r)}T_{r-\tau(r)}^{-1}(x)=x_{1}-x_{2},$
where $x_{1}=T_{r}T_{r-\tau(r)}^{-1}(x)\in\mathrm{Im}(T_{r}),~x_{2}=T_{\tau(r)}T_{r-\tau(r)}^{-1}(x)\in\mathrm{Im}(T_{\tau(r)})$. The proof is finished. ∎
Proposition 4.12.
Let $(A,\succ_{A},\prec_{A},\ast_{A},\Delta_{\succ,r},\Delta_{\prec,r},\delta_{r})$ be a factorizable noncommutative pre-Poisson bialgebra. Then the double noncommutative pre-Poisson algebra $(D=A\oplus A^{*},\succ_{D},\prec_{D},\ast_{D})$ is isomorphic to the direct sum $A\oplus A$ of noncommutative pre-Poisson algebras, where $\succ_{D},\prec_{D}$ and $\ast_{D}$ are given by Eqs. (92)-(94) respectively.
Proof.
Since $(A,\succ_{A},\prec_{A},\ast_{A},\Delta_{\succ,r},\Delta_{\prec,r},\delta_{r})$ is a factorizable noncommutative pre-Poisson bialgebra, $T_{r-\tau(r)}$ is a linear isomorphism and $r-\tau(r)$ is invariant. Define $\varphi:A\oplus A^{*}\longrightarrow A\oplus A$ by
(126) $\varphi(x,\zeta)=(x+T_{r}(\zeta),x+T_{\tau(r)}(\zeta)),~\forall~(x,\zeta)\in A\oplus A^{*}.$
Then $\varphi$ is a bijection. By Eq. (124), we get
$\langle\eta,-L_{\prec_{A^{*}}}^{*}(\zeta)x+T_{r}(R_{\cdot_{A}}^{*}(x)\zeta\rangle$
$=$ $\langle-\zeta\prec\eta,x\rangle+\langle\eta,T_{r}(R_{\cdot_{A}}^{*}(x)\zeta\rangle$
$=$ $\langle R_{\succ}^{*}(T_{r}(\zeta))\eta-L_{\cdot}^{*}(T_{\tau(r)}(\eta))\zeta,x\rangle+\langle\eta,T_{r}(R_{\cdot_{A}}^{*}(x)\zeta\rangle$
$=$ $\langle x\succ_{A}T_{r}(\zeta),\eta\rangle-\langle\zeta,T_{\tau(r)}(\eta)\cdot_{A}x\rangle+\langle\eta,T_{r}(R_{\cdot_{A}}^{*}(x)\zeta\rangle$
$=$ $\langle x\succ_{A}T_{r}(\zeta),\eta\rangle-\langle T_{r}(R_{\cdot_{A}}^{*}(x)\zeta),\eta\rangle+\langle\eta,T_{r}(R_{\cdot_{A}}^{*}(x)\zeta\rangle$
$=$ $\langle x\succ_{A}T_{r}(\zeta),\eta\rangle,$
which implies that
(127) $x\succ_{A}T_{r}(\zeta)=-L_{\prec_{A^{*}}}^{*}(\zeta)x+T_{r}(R_{\cdot_{A}}^{*}(x)\zeta.$
Similarly, using Eqs. (124) and (118), we have
(128) $x\succ_{A}T_{\tau(r)}(\zeta)=-L_{\prec_{A^{*}}}^{*}(\zeta)x+T_{\tau(r)}(R_{\cdot_{A}}^{*}(x)\zeta.$
By Eqs. (93) and (127)-(128), we get
$\varphi(x\succ_{D}\zeta)=\varphi(-L_{\prec_{A^{*}}}^{*}(\zeta)x+R_{\cdot_{A}}^{*}(x)\zeta)$
$=$ $(-L_{\prec_{A^{*}}}^{*}(\zeta)x+T_{r}(R_{\cdot_{A}}^{*}(x)\zeta),-L_{\prec_{A^{*}}}^{*}(\zeta)x+T_{\tau(r)}(R_{\cdot_{A}}^{*}(x)\zeta))$
$=$ $(x\succ_{A}T_{r}(\zeta),x\succ_{A}T_{\tau(r)}(\zeta))=(x,x)\succ(T_{r}(\zeta),T_{\tau(r)}(\zeta))$
$=$ $\varphi(x)\succ_{A}\varphi(\zeta).$
Analogously, $\varphi(\zeta\succ_{D}x)=\varphi(\zeta)\succ_{A}\varphi(x)$. Thus, $\varphi((x,\zeta)\succ_{D}(y,\eta))=\varphi(x,\zeta)\succ\varphi(y,\eta)$ for all $(x,\zeta),(y,\eta)\in A\oplus A^{*}$. Analogously, $\varphi((x,\zeta)\prec_{D}(y,\eta))=\varphi(x,\zeta)\prec\varphi(y,\eta)$ and $\varphi((x,\zeta)\ast_{D}(y,\eta))=\varphi(x,\zeta)\ast\varphi(y,\eta)$. In all, $\varphi$ is an isomorphism of noncommutative pre-Poisson algebras. The proof is completed. ∎
Theorem 4.13.
Let $(A,\succ,\prec,\ast,\Delta_{\succ,r},\Delta_{\prec,r},\delta_{r})$ be a noncommutative pre-Poisson bialgebra. Assume that $\{e_{1},...,e_{n}\}$ is a basis of $A$ and $\{e^{*}_{1},...,e^{*}_{n}\}$ is the dual basis. Let
$r=\sum_{i=1}^{n}e_{i}\otimes e_{i}^{*}\in A\otimes A^{*}\subseteq D\otimes D.$
Then $(D,\succ_{D},\prec_{D},\ast_{D},\Delta_{\succ,r},\Delta_{\prec,r},\delta_{r})$ is a factorizable noncommutative pre-Poisson bialgebra with $\Delta_{\succ,r},\Delta_{\prec,r},\delta_{r}$ given by Eqs. (67)-(68) and (78) respectively.
Proof.
On the basis of Theorem 2.21 [30] and Theorem 2.18 [29], we get that $r-\tau(r)$ is invariant and $S(r)=D(r)=0$. Thus, $(A,\succ,\prec,\ast,\Delta_{\succ,r},\Delta_{\prec,r},\delta_{r})$ is a quasi-triangular noncommutative pre-Poisson bialgebra. Furthermore, the linear maps $T_{r},T_{\tau(r)}:D^{*}\longrightarrow D$ are respectively defined by $T_{r}(\zeta,x)=\zeta,~T_{\tau(r)}(\zeta,x)=x$ for all $x\in A,\zeta\in A^{*}$. Thus, $T_{r-\tau(r)}(\zeta,x)=(\zeta,-x)$ is a linear isomorphism. Hence, $(D,\succ_{D},\prec_{D},\ast_{D},\Delta_{\succ,r},\Delta_{\prec,r},\delta_{r})$ is a factorizable noncommutative pre-Poisson bialgebra. ∎
5. Relative Rota-Baxter operators and quadratic Rota-Baxter noncommutative pre-Poisson algebras
5.1. Relative Rota-Baxter operators and the NPP-YBE
This section focuses on solutions of the NPP-YBE with invariant skew-symmetric parts, studying them specifically via relative Rota-Baxter operators of weight $\lambda$ on noncommutative pre-Poisson algebras
Proposition 5.1.
Let $(A,\succ,\prec,\ast)$ be a coherent noncommutative pre-Poisson algebra and $r\in A\otimes A$ be skew-symmetric and invariant. Define the operations $\succ_{r},\prec_{r},\ast_{r}:A^{*}\otimes A^{*}\longrightarrow A^{*}$ by
(129) $\zeta\succ_{r}\eta=R_{\cdot}^{*}(T_{r}(\zeta))\eta=-L_{\prec}^{*}(T_{r}(\eta))\zeta,$
(130) $\zeta\prec_{r}\eta=-R_{\succ}^{*}(T_{r}(\zeta))\eta=L_{\cdot}^{*}(T_{r}(\eta))\zeta,$
(131) $\zeta\ast_{r}\eta=-\mathrm{ad}^{*}(T_{r}(\zeta))\eta=R_{\ast}^{*}(T_{r}(\eta))\zeta,~~\forall~\zeta,\eta\in A^{*}.$
Then $(A^{*},\succ_{r},\prec_{r},\ast_{r},R_{\cdot}^{*},-L_{\prec}^{*},-R_{\succ}^{*},L_{\cdot}^{*},R^{*}_{\ast}-L_{\ast}^{*},R^{*}_{\ast})$ is an A-noncommutative pre-Poisson algebra and $(A^{*},\cdot_{r},[\ ,\ ]_{r},R_{\prec}^{*},L_{\succ}^{*},-L_{\ast}^{*})$ is an A-noncommutative Poisson algebra, where the Poisson algebra structure $(\cdot_{r},[\ ,\ ]_{r})$ on $A^{*}$ is given by
(132) $\zeta\cdot_{r}\eta=R_{\cdot}^{*}(T_{r}(\zeta))\eta-R_{\succ}^{*}(T_{r}(\zeta))\eta=-L_{\prec}^{*}(T_{r}(\eta))\zeta+L_{\cdot}^{*}(T_{r}(\eta))\zeta,$
(133) $[\zeta,\eta]_{r}=\mathrm{ad}^{*}(T_{r}(\eta))\zeta-\mathrm{ad}^{*}(T_{r}(\zeta))\eta=R_{\ast}^{*}(T_{r}(\zeta))\eta-R_{\ast}^{*}(T_{r}(\eta))\zeta.$
Proof.
By Eqs. (129), (131), (133), (117), (120) and (6), we obtain
$\zeta\ast_{r}L_{\prec}^{*}(x)\eta+\eta\succ_{r}(R_{\ast}^{*}(x)\zeta)-L_{\prec}^{*}(x)[\zeta,\eta]_{r}$
$=$ $-\mathrm{ad}^{*}(T_{r}(\zeta))L_{\prec}^{*}(x)\eta+R_{\cdot}^{*}(T_{r}(\eta))R_{\ast}^{*}(x)\zeta+L_{\prec}^{*}(x)(\mathrm{ad}^{*}(T_{r}(\zeta))\eta-\mathrm{ad}^{*}(T_{r}(\eta))\zeta)$
$=$ $R_{\ast}^{*}(T_{r}(L_{\prec}^{*}(x)\eta))\zeta+R_{\cdot}^{*}(T_{r}(\eta))R_{\ast}^{*}(x)\zeta-L_{\prec}^{*}(x)(R_{\ast}^{*}(T_{r}(\eta))\zeta+\mathrm{ad}^{*}(T_{r}(\eta))\zeta)$
$=$ $-R_{\ast}^{*}(R_{\succ}(x)T_{r}(\eta))\zeta+R_{\cdot}^{*}(T_{r}(\eta))R_{\ast}^{*}(x)\zeta-L_{\prec}^{*}(x)L_{\ast}^{*}(T_{r}(\eta))\zeta$
$=$ $0,$
which means that Eq. (46) holds in the case that $l_{\ast}=R_{\ast}^{*}-L_{\ast}^{*},~r_{\ast}=R_{\ast}^{*},~l_{\succ}=R_{\cdot}^{*},~r_{\succ}=-L_{\prec}^{*},~l_{\prec}=-R_{\succ}^{*},~r_{\prec}=L_{\cdot}^{*}$. Similarly, we prove that Eqs. (46)-(54) hold and $(A^{*},\ast_{r},R^{*}_{\ast}-L_{\ast}^{*},R^{*}_{\ast})$ is an A-pre-Lie algebra. By Lemma 3.4 [30], $(A^{*},\succ_{r},\prec_{r},R_{\cdot}^{*},-L_{\prec}^{*},-R_{\succ}^{*},L_{\cdot}^{*})$ is an A-dendriform algebra. Thus, $(A^{*},\succ_{r},\prec_{r},\ast_{r},R_{\cdot}^{*},-L_{\prec}^{*},-R_{\succ}^{*},L_{\cdot}^{*},R^{*}_{\ast}-L_{\ast}^{*},R^{*}_{\ast})$ is an A-noncommutative pre-Poisson algebra. Then $(A^{*},\cdot_{r},[\ ,\ ]_{r},R_{\prec}^{*},L_{\succ}^{*},-L_{\ast}^{*})$ is an A-noncommutative Poisson algebra holds naturally. This completes the proof. ∎
Proposition 5.2.
Let $(A,\succ,\prec,\ast)$ be a noncommutative pre-Poisson algebra and $r\in A\otimes A$. Then
  1. (a)
    $r$ is a solution of the equation $D_{2}(r)=r_{13}\cdot r_{23}-r_{12}\succ r_{13}-r_{23}\prec r_{12}=0$ in $(A,\succ,\prec)$ if and only if the following equation holds:
    $T_{r}(\zeta)\cdot T_{r}(\eta)=T_{r}(R_{\prec}^{*}(T_{r}(\zeta))\eta+L_{\succ}^{*}(T_{\tau(r)}(\eta))\zeta),~~\forall~\zeta,\eta\in A^{*}.$
  2. (b)
    $r$ is a solution of the D-equation $D(r)=r_{12}\cdot r_{13}-r_{23}\succ r_{12}-r_{13}\prec r_{23}=0$ in $(A,\succ,\prec)$ if and only if the following equation holds:
    $T_{r}(\zeta)\prec T_{r}(\eta)=T_{r}(L_{\cdot}^{*}(T_{\tau(r)}(\eta))\zeta-R_{\succ}^{*}(T_{r}(\zeta))\eta),~~\forall~\zeta,\eta\in A^{*}.$
  3. (c)
    $r$ is a solution of the equation $D_{3}(r)=r_{23}\cdot r_{12}-r_{13}\succ r_{23}-r_{12}\prec r_{13}=0$ if and only if the following equation holds:
    $T_{r}(\zeta)\succ T_{r}(\eta)=T_{r}(R_{\cdot}^{*}(T_{r}(\zeta))\eta-L_{\prec}^{*}(T_{\tau(r)}(\eta))\zeta),~~\forall~\zeta,\eta\in A^{*}.$
  4. (d)
    $r$ is a solution of the S-equation $S(r)=r_{12}\ast r_{23}-r_{12}\ast r_{13}+[r_{13},r_{23}]=0$ if and only if the following equation holds:
    $T_{r}(\zeta),T_{r}(\eta)]=T_{r}(L_{\ast}^{*}(T_{\tau(r)}(\eta))\zeta-L_{\ast}^{*}(T_{r}(\zeta))\eta),~~\forall~\zeta,\eta\in A^{*}.$
  5. (e)
    $r$ is a solution of the equation $S_{1}(r)=r_{23}\ast r_{12}-r_{23}\ast r_{13}-[r_{12},r_{13}]=0$ if and only if the following equation holds:
    $T_{r}(\zeta)\ast T_{r}(\eta)=T_{r}(R_{\ast}^{*}(T_{r}(\eta))\zeta-\mathrm{ad}^{*}(T_{\tau(r)}(\zeta))\eta),~~\forall~\zeta,\eta\in A^{*}.$
Proof.
According to Eq. (107), we have for all $\zeta,\eta,\theta\in A^{*}$,
$\langle\zeta\otimes\eta\otimes\theta,r_{13}\cdot r_{23}\rangle$ $=\sum_{i,j}\langle\zeta\otimes\eta\otimes\theta,a_{i}\otimes a_{j}\otimes b_{i}\cdot b_{j}\rangle$
$=\sum_{i,j}\langle\theta,b_{i}\cdot b_{j}\rangle\langle\zeta,a_{i}\rangle\langle\eta,a_{j}\rangle=\langle\theta,T_{r}(\zeta)\cdot T_{r}(\eta)\rangle,$
$\langle\zeta\otimes\eta\otimes\theta,r_{23}\prec r_{12}\rangle$ $=\sum_{i,j}\langle\zeta\otimes\eta\otimes\theta,a_{j}\otimes a_{i}\prec b_{j}\otimes b_{i}\rangle$
$=\sum_{i,j}\langle\zeta,a_{j}\rangle\langle\theta,b_{i}\rangle\langle\eta,b_{i}\prec b_{j}\rangle=\langle\eta,T_{r}(\zeta)\prec T_{\tau(r)}(\theta)\rangle$
$=\langle L_{\prec}^{*}(T_{r}(\zeta))\eta,T_{\tau(r)}(\theta)\ \rangle=\langle\theta,T_{r}(L_{\prec}^{*}(T_{r}(\zeta))\eta)\rangle,$
$\langle\zeta\otimes\eta\otimes\theta,r_{12}\succ r_{13}\rangle$ $=\sum_{i,j}\langle\zeta\otimes\eta\otimes\theta,a_{i}\succ a_{j}\otimes b_{i}\otimes b_{j}\rangle$
$=\sum_{i,j}\langle\eta,b_{i}\rangle\langle\zeta,a_{i}\succ a_{j}\rangle\langle\theta,b_{j}\rangle=\langle\zeta,T_{\tau(r)}(\eta)\succ T_{\tau(r)}(\theta)\rangle$
$=\langle L_{\succ}^{*}(T_{\tau(r)}(\eta))\zeta,T_{\tau(r)}(\theta)\rangle=\langle\theta,T_{r}(L_{\succ}^{*}(T_{\tau(r)}(\eta))\zeta\rangle.$
Thus, Item (a) holds. Similarly, Items (b)-(e) hold.∎
Proposition 5.3.
Let $(A,\succ,\prec,\ast)$ be a noncommutative pre-Poisson algebra and $r\in A\otimes A$. Assume that $r-\tau(r)$ is invariant. Then
  1. (a)
    $r$ is a solution of the S-equation $S(r)=0$ if and only if $r$ is a solution of $S_{1}(r)=0$.
  2. (b)
    $r$ is a solution of $D(r)=0$ if and only if $r$ is a solution of $D_{3}(r)=D_{2}(r)=0$.
Proof.
Note that $D_{3}(r)+D_{2}(r)=-D(r)$ and
$S_{1}(r)=$ $-\sigma_{123}S(r)+r_{23}\ast(r_{31}-r_{13})+[r_{21},r_{31}-r_{13}]+r_{23}\ast(r_{12}-r_{21})+[r_{21}-r_{12},r_{13}]$
$=$ $-\sigma_{123}S(r)-\sum_{i}(I\otimes I\otimes L_{\ast}(b_{i})+\mathrm{ad}(b_{i})\otimes I\otimes I)(\tau\otimes I)(a_{i}\otimes(r-\tau(r)))$
$+(I\otimes L_{\ast}(a_{i})+\mathrm{ad}(a_{i})\otimes I)(r-\tau(r))\otimes b_{i},$
$D_{3}(r)=$ $\sigma_{123}D(r)+r_{23}\cdot(r_{12}-r_{21})-(r_{12}-r_{21})\prec r_{13}-(r_{13}-r_{31})\succ r_{23}-r_{21}\prec(r_{13}-r_{31})$
$=$ $\sigma_{123}D(r)+\sum_{i}(I\otimes L_{\cdot}(a_{i})-R_{\prec}(a_{i})\otimes I)(r-\tau(r))\otimes b_{i}$
$-(I\otimes I\otimes R_{\succ}(b_{i})+L_{\prec}(b_{i})\otimes I\otimes I)(\tau\otimes I)(a_{i}\otimes(r-\tau(r))),$
$D_{2}(r)=$ $\sigma_{132}D(r)+r_{13}\cdot(r_{23}-r_{32})-(r_{23}-r_{32})\prec r_{12}+(r_{13}-r_{31})\cdot r_{32}+r_{12}\succ(r_{31}-r_{13})$
$=$ $\sigma_{132}D(r)+\sum_{i}a_{i}\otimes(I\otimes L_{\cdot}(b_{i})-R_{\prec}(b_{i})\otimes I)(r-\tau(r))$
$+(I\otimes I\otimes R_{\cdot}(a_{i})-L_{\succ}(a_{i})\otimes I\otimes I)(\tau\otimes I)(b_{i}\otimes(r-\tau(r))).$
Thus, the conclusions hold. ∎
The following Theorem is a generalization of Theorem 3.24.
Theorem 5.4.
Let $(A,\succ,\prec,\ast)$ be a coherent noncommutative pre-Poisson algebra and $r\in A\otimes A$. Assume that $r-\tau(r)$ is invariant. Then the following conditions are equivalent.
  1. (a)
    $r$ is a solution of the NPP-YBE in $(A,\succ,\prec,\ast)$ such that $(A,\succ,\prec,\ast,\Delta_{\succ,r},\Delta_{\prec,r},\delta_{r})$ is a quasi-triangular noncommutative pre-Poisson bialgebra with $\Delta_{\succ,r},\Delta_{\prec,r},\delta_{r}$ given by Eqs. (67)-(68) and (78) respectively.
  2. (b)
    $T_{r}$ is a relative Rota-Baxter operator of weight 1-1 on $(A,\succ,\prec,\ast)$ with respect to the A-noncommutative pre-Poisson algebra $(A^{*},\succ_{r-\tau(r)},\prec_{r-\tau(r)},\ast_{r-\tau(r)},R_{\cdot}^{*},-L_{\prec}^{*},-R_{\succ}^{*},L_{\cdot}^{*},R^{*}_{\ast}-L_{\ast}^{*},R^{*}_{\ast})$, that is,
    (134) $T_{r}(\zeta)\succ T_{r}(\eta)=T_{r}(R_{\cdot}^{*}(T_{r}(\zeta))\eta-L_{\prec}^{*}(T_{r}(\eta))\zeta-\zeta\succ_{r-\tau(r)}\eta),$
    (135) $T_{r}(\zeta)\prec T_{r}(\eta)=T_{r}(-R_{\succ}^{*}(T_{r}(\zeta))\eta+L_{\cdot}^{*}(T_{r}(\eta))\zeta-\zeta\prec_{r-\tau(r)}\eta),$
    (136) $T_{r}(\zeta)\ast T_{r}(\eta)=T_{r}((R^{*}_{\ast}-L_{\ast}^{*})(T_{r}(\zeta))\eta+R_{\ast}^{*}(T_{r}(\eta))\zeta-\zeta\ast_{r-\tau(r)}\eta).$
  3. (c)
    $T_{r}$ is a relative Rota-Baxter operator of weight 1-1 on $(A,\cdot,[\ ,\ ])$ with respect to the A-noncommutative Poisson algebra $(A^{*},\cdot_{r-\tau(r)},[\ ,\ ]_{r-\tau(r)},R_{\prec}^{*},L_{\succ}^{*},-L_{\ast}^{*})$, that is,
    (137) $T_{r}(\zeta)\cdot T_{r}(\eta)=T_{r}(R_{\prec}^{*}(T_{r}(\zeta))\eta+L_{\succ}^{*}(T_{r}(\eta))\zeta-\zeta\cdot_{r-\tau(r)}\eta),$
    (138) $[T_{r}(\zeta),T_{r}(\eta)]=T_{r}(L_{\ast}^{*}(T_{r}(\eta))\zeta-L_{\ast}^{*}(T_{r}(\zeta))\eta-[\zeta,\eta]_{r-\tau(r)}),$
for all $\zeta,\eta\in A^{*}$, where $\succ_{r-\tau(r)},\prec_{r-\tau(r)},\ast_{r-\tau(r)}$ and $\cdot_{r-\tau(r)},[\ ,\ ]_{r-\tau(r)}$ are given by Eqs. (129)-(133).
Proof.
Based on Proposition 5.2 and Proposition 5.3, we can check the statement holds. ∎
5.2. Quadratic Rota-Baxter noncommutative pre-Poisson algebras and factorizable noncommutative pre-Poisson bialgebras
In this section, we first introduce the notion of quadratic Rota-Baxter noncommutative pre-Poisson algebras. We then establish the relationship between factorizable noncommutative pre-Poisson bialgebras and these quadratic Rota-Baxter noncommutative pre-Poisson algebras.
Definition 5.5.
Assume that $(A,\succ,\prec,\ast,P)$ is a Rota-Baxter noncommutative pre-Poisson algebra of weight $\lambda$ and $(A,\succ,\prec,\ast,\omega)$ a quadratic noncommutative pre-Poisson algebra. Then $(A,\succ,\prec,\ast,P,\omega)$ is called a quadratic Rota-Baxter noncommutative pre-Poisson algebra of weight $\lambda$ if the following condition holds:
(139) $\omega(P(x),y)+\omega(x,P(y))+\lambda\omega(x,y)=0,~\forall~x,y\in A.$
Definition 5.6.
Let $(A,\cdot,[\ ,\ ],P)$ be a Rota-Baxter noncommutative Poisson algebra of weight $\lambda$ and $(A,\cdot,[\ ,\ ],\omega)$ be a symplectic noncommutative Poisson algebra if the following condition holds:
(140) $\omega(P(x),y)+\omega(x,P(y))+\lambda\omega(x,y)=0,~\forall~x,y\in A.$
Then $(A,\cdot,[\ ,\ ],P,\omega)$ is called a symplectic Rota-Baxter noncommutative Poisson algebra of weight $\lambda$.
Observe that for all $x,y\in A$,
$\lambda\omega(x,y)+\omega(-\lambda(x)-P(x),y)+\omega(x,-\lambda(y)-P(y))$
$=$ $-\lambda\omega(x,y)-\omega(P(x),y)-\omega(x,P(y)),$
from which we obtain the following conclusions.
Proposition 5.7.
Let $(A,\succ,\prec,\ast,P,\omega)$ be a quadratic Rota-Baxter noncommutative pre-Poisson algebra and let $P:A\longrightarrow A$ be a linear map. Then $(A,\succ,\prec,\ast,P,\omega)$ is a quadratic Rota-Baxter noncommutative pre-Poisson algebra of weight $\lambda$ if and only if $(A,\succ,\prec,\ast,-\lambda I-P,\omega)$ is a quadratic Rota-Baxter noncommutative pre-Poisson algebra of weight $\lambda$.
Proposition 5.8.
Let $(A,\cdot,[\ ,\ ],\omega)$ be a quadratic noncommutative Poisson algebra and let $P:A\longrightarrow A$ be a linear map. Then $(A,\cdot,[\ ,\ ],P,\omega)$ is a symplectic Rota-Baxter noncommutative Poisson algebra of weight $\lambda$ if and only if $(A,\cdot,[\ ,\ ],-\lambda I-P,\omega)$ is a symplectic Rota-Baxter noncommutative Poisson algebra of weight $\lambda$.
Based on Proposition 2.15, the following can be obtained easily.
Theorem 5.9.
Suppose that $(A,\cdot,[\ ,\ ],P,\omega)$ is a symplectic Rota-Baxter noncommutative Poisson algebra of weight $\lambda$. Then $(A,\succ,\prec,\ast,P,\omega)$ is a quadratic Rota-Baxter noncommutative pre-Poisson algebra of weight $\lambda$, where $\succ,\prec,\ast$ are defined by Eqs. (64) and (65) respectively. On the other hand, let $(A,\succ,\prec,\ast,P,\omega)$ be a quadratic Rota-Baxter noncommutative pre-Poisson algebra of weight $\lambda$. Then $(A,\cdot,[\ ,\ ],P,\omega)$ is a symplectic Rota-Baxter noncommutative Poisson algebra of weight $\lambda$, where $x\cdot y=x\succ y+x\prec y,~[x,y]=x\ast y-y\ast x$.
Let $\omega$ be a non-degenerate bilinear form on a vector space $A$. Then there is an isomorphism $\omega^{\sharp}:A\longrightarrow A^{*}$ given by
(141) $\omega(x,y)=\langle\omega^{\sharp}(x),y\rangle,~~\forall~x,y\in A.$
Define an element $r_{\omega}\in A\otimes A$ such that $T_{r_{\omega}}=(\omega^{\sharp})^{-1}$, that is,
(142) $\langle T_{r_{\omega}}(\zeta),\eta\rangle=\langle r_{\omega},\zeta\otimes\eta\rangle=\langle(\omega^{\sharp})^{-1}(\zeta),\eta\rangle,\ \ \forall~\zeta,\eta\in A^{*}.$
Lemma 5.10.
Let $(A,\succ,\prec,\ast)$ be a noncommutative pre-Poisson algebra and $\omega$ be a non-degenerate bilinear form on $A$. Then $(A,\succ,\prec,\ast,\omega)$ is a quadratic noncommutative pre-Poisson algebra if and only if the corresponding $r_{\omega}\in A\otimes A$ given by Eq. (142) is skew-symmetric and invariant.
Proof.
It is obvious that $\omega$ is skew-symmetric if and only if $r_{\omega}$ is skew-symmetric. For all $x,y\in A$, put $\omega^{\sharp}(x)=\zeta,\omega^{\sharp}(y)=\eta,\omega^{\sharp}(z)=\theta$ with $\zeta,\eta,\theta\in A^{*}$. If $\omega$ is skew-symmetric, we have
$\omega(x\ast y,z)+\omega(y,[x,z])=$ $\langle-\omega^{\sharp}(z),x\ast(\omega^{\sharp})^{-1}(\eta)\rangle+\langle\omega^{\sharp}(y),[x,(\omega^{\sharp})^{-1}(\theta)]\rangle$
$=$ $-\langle\theta,x\ast T_{r_{\omega}}(\eta)\rangle+\langle\eta,[x,T_{r_{\omega}}(\theta)]\rangle$
$=$ $\langle-R_{\ast}^{*}(T_{r_{\omega}}(\eta))\theta-\mathrm{ad}^{*}(T_{r_{\omega}}(\theta))\eta,x\rangle.$
Thus, if $\omega$ is skew-symmetric, $\omega(x\ast y,z)=\omega(y,[x,z])$ if and only if $R_{\ast}^{*}(T_{r_{\omega}}(\eta))\theta=-\mathrm{ad}^{*}(T_{r_{\omega}}(\theta))\eta$. By the same token, if $\omega$ is skew-symmetric, $\omega(x\succ y,z)=\omega(y,z\cdot x)$ if and only if $L_{\cdot}^{*}(T_{r_{\omega}}(\zeta))\eta=-R_{\succ}^{*}(T_{r_{\omega}}(\eta))\zeta).$ Combining Proposition 4.3, we complete the proof. ∎
Proposition 5.11.
Let $(A,\succ,\prec,\ast,\omega)$ be a quadratic noncommutative pre-Poisson algebra and $r\in A\otimes A$. Assume that $r-\tau(r)$ is invariant. Define a linear map
(143) $P:A\longrightarrow A,\ \ \ P(x)=-T_{r}\omega^{\sharp}(x),~~\forall~x\in A.$
Then $r$ is a solution of the NPP-YBE in $(A,\succ,\prec,\ast)$ if and only if $P$ satisfies
(144) $P(x)\succ P(y)=P(P(x)\succ y+x\succ P(y)-x\succ T_{r-\tau(r)}\omega^{\sharp}(y)),$
(145) $P(x)\prec P(y)=P(P(x)\prec y+x\prec P(y)-x\prec T_{r-\tau(r)}\omega^{\sharp}(y))$
(146) $P(x)\ast P(y)=P(P(x)\ast y+x\ast P(y)-x\ast T_{r-\tau(r)}\omega^{\sharp}(y)),~~\forall~x,y\in A.$
Proof.
By Lemma 5.10, $r_{\omega}$ is skew-symmetric and invariant. In light of Proposition 4.3, Eqs. (118)-(120) and (130), for all $x,y\in A$, put $\omega^{\sharp}(x)=\zeta,\omega^{\sharp}(y)=\eta$ with $\zeta,\eta\in A^{*}$, we get
$P(x)\succ P(y)=T_{r}\omega^{\sharp}(x)\succ T_{r}\omega^{\sharp}(y)=T_{r}(\zeta)\succ T_{r}(\eta),$
$P(P(x)\succ y)=T_{r}\omega^{\sharp}(T_{r}(\zeta)\succ T_{r_{\omega}}(\eta))=T_{r}\omega^{\sharp}T_{r_{\omega}}(R_{\cdot}^{*}(T_{r}(\zeta))\eta)=T_{r}(R_{\cdot}^{*}(T_{r}(\zeta))\eta),$
$P(x\succ P(y))=T_{r}\omega^{\sharp}(T_{r_{\omega}}(\zeta)\succ T_{r}(\eta))=-T_{r}\omega^{\sharp}T_{r_{\omega}}(L_{\prec}^{*}(T_{r}(\eta))\zeta)=-T_{r}(L_{\prec}^{*}(T_{r}(\eta))\zeta),$
$P(x\succ T_{r-\tau(r)}\omega^{\sharp}(y))=-T_{r}\omega^{\sharp}(T_{r_{\omega}}(\zeta)\succ T_{r-\tau(r)}(\eta))=T_{r}\omega^{\sharp}T_{r_{\omega}}(L_{\prec}^{*}(T_{r-\tau(r)}(\eta))\zeta)$
$=$ $T_{r}((L_{\prec}^{*}(T_{r+\tau(r)}(\eta))\zeta)=-T_{r}(\zeta\succ_{r-\tau(r)}\eta).$
Thus, Eq. (144) holds if and only if Eq. (135) holds. By the same token, we can verify that Eq. (145)-(147) hold if and only if Eqs. (134)-(135) hold. Combining Theorem 5.4, we get the conclusion.
Lemma 5.12.
Let $A$ be a vector space and $\omega$ be a non-degenerate skew-symmetric bilinear form. Let $r\in A\otimes A$, $\lambda\in k$ and a linear map $P$ given by Eq. (143). Then $r$ satisfies
(147) $r-\tau(r)=\lambda r_{\omega}$
if and only if $P$ satisfies Eq. (139).
Proof.
For all $x,y\in A$, put $\omega^{\sharp}(x)=\zeta,\omega^{\sharp}(y)=\eta$ with $\zeta,\eta\in A^{*}$.
$\omega(P(x),y)=-\omega(y,P(x))=\langle\omega^{\sharp}(y),T_{r}\omega^{\sharp}(x)\rangle=\langle\eta,T_{r}(\zeta)\rangle=\langle r,\zeta\otimes\eta\rangle,$
$\omega(x,P(y))=-\langle\omega^{\sharp}(x),T_{r}\omega^{\sharp}(y)\rangle=-\langle\zeta,T_{r}(\eta)\rangle=-\langle r,\eta\otimes\zeta\rangle=-\langle\tau(r),\zeta\otimes\eta\rangle,$
$\lambda\omega(x,y)=-\lambda\omega(y,x)=-\lambda\langle\omega^{\sharp}(y),(\omega^{\sharp})^{-1}\omega^{\sharp}(x)\rangle=-\lambda\langle r_{\omega},\zeta\otimes\eta\rangle.$
Thus, Eq. (147) holds if and only if Eq. (139) holds. ∎
Corollary 5.13.
Suppose that $(A,\succ,\prec,\ast,P,\omega)$ is a quadratic Rota-Baxter noncommutative pre-Poisson algebra of weight 0. Then there is a triangular noncommutative pre-Poisson bialgebra $(A,\succ,\prec,\ast,\Delta_{\succ,r},\Delta_{\prec,r},\delta_{r})$ with $\Delta_{\succ,r},\Delta_{\prec,r},\delta_{r}$ given by Eqs. (67)-(68) and (78), where $r\in A\otimes A$ given by $T_{r}(\zeta)=P(\omega^{\sharp})^{-1}(\zeta)$ for all $\zeta\in A^{*}$.
Proof.
Since $\tau(r)=\lambda r_{\omega}=0, rr$ is symmetric. On the basis of Proposition 5.11 and Lemma 5.12, we get the conclusion. ∎
Theorem 5.14.
Let $(A,\succ,\prec,\ast,\Delta_{\succ,r},\Delta_{\prec,r},\delta_{r})$ be a factorizable noncommutative pre-Poisson bialgebra with $r\in A\otimes A$. Then $(A,\succ,\prec,\ast,P,\omega)$ is a quadratic Rota-Baxter noncommutative pre-Poisson algebra of weight $\lambda$ with $P$ given by Eq. (143), and $\omega$ is given by
(148) $\omega(x,y)=\lambda\langle T_{r-\tau(r)}^{-1}(x),y\rangle,~~\forall~x,y\in A.$
Conversely, let $(A,\succ,\prec,\ast,P,\omega)$ be a quadratic Rota-Baxter noncommutative pre-Poisson algebra of weight $\lambda~(\lambda\neq 0)$. Then there exists a factorizable noncommutative pre-Poisson bialgebra $(A,\succ,\prec,\ast,\Delta_{\succ,r},\Delta_{\prec,r},\delta_{r})$ with $\Delta_{\succ,r},\Delta_{\prec,r},\delta_{r}$ defined by Eqs. (67)-(68) and (78) respectively, where $r\in A\otimes A$ is given through the operator form $T_{r}=-P(\omega^{\sharp})^{-1}$.
Proof.
On the one hand, since $(A,\succ,\prec,\ast,\Delta_{\succ,r},\Delta_{\prec,r},\delta_{r})$ is a factorizable noncommutative pre-Poisson bialgebra, $r-\tau(r)$ is invariant and $T_{r-\tau(r)}$ is a linear isomorphism. By Proposition 5.11 and Lemma 5.12, we get that $(A,\succ,\prec,\ast,P,\omega)$ is a quadratic Rota-Baxter noncommutative pre-Poisson algebra of weight $\lambda$, where $\omega^{\sharp}=\lambda T_{r-\tau(r)}^{-1}$. Conversely, assume that $(A,\succ,\prec,\ast,P,\omega)$ is a quadratic Rota-Baxter noncommutative pre-Poisson algebra of weight $\lambda~(\lambda\neq 0)$. In light of Lemma 5.10, Lemma 5.12 and Proposition 5.11, $r-\tau(r)$ is invariant, $T_{r-\tau(r)}=\lambda(\omega^{\sharp})^{-1}$ is a linear isomorphism and $r$ is a solution of the NPP-YBE in $(A,\succ,\prec,\ast)$. Thus, $(A,\succ,\prec,\ast,\Delta_{\succ,r},\Delta_{\prec,r},\delta_{r})$ is a factorizable noncommutative pre-Poisson bialgebra. ∎
By Theorem 5.9 and Theorem 5.14, we have the following correspondence:
$(A,\cdot,[\ ,\ ],P,\omega){\overset{\text{Theorem 5.10}}{\longleftrightarrow}}(A,\succ,\prec,\ast,P,\omega){\overset{\text{Theorem 5.15}}{\longleftrightarrow}}(A,\succ,\prec,\ast,\Delta_{\succ,r},\Delta_{\prec,r},\delta_{r}).$
According to Proposition 5.7 and Theorem 5.14, a quadratic Rota-Baxter noncommutative pre-Poisson algebra $(A,\succ,\prec,\ast,P,\omega)$ of a non-zero weight yields a factorizable noncommutative pre-Poisson bialgebra. Consequently, $(A,\succ,\prec,\ast,-\lambda I-P,\omega)$ yields one as well. Furthermore, Proposition 4.10 and Theorem 5.14 imply that if $(A,\succ,\prec,\ast,\Delta_{\succ,r},\Delta_{\prec,r},\delta_{r})$ is such a factorizable bialgebra, then $(A,\succ,\prec,\ast,\Delta_{\succ,\tau(r)},\Delta_{\prec,\tau(r)},\delta_{\tau(r)})$ also produces a quadratic Rota-Baxter noncommutative pre-Poisson algebra of weight $\lambda$. Indeed, we have
Proposition 5.15.
Assume that $(A,\succ,\prec,\ast,\Delta_{\succ,r},\Delta_{\prec,r},\delta_{r})$ is a factorizable noncommutative pre-Poisson bialgebra which corresponds to a quadratic Rota-Baxter noncommutative pre-Poisson algebra of non-zero weight $\lambda$. Then the factorizable noncommutative pre-Poisson bialgebra $(A,\succ,\prec,\ast,\Delta_{\succ,\tau(r)},\Delta_{\prec,\tau(r)},\delta_{\tau(r)})$ corresponds to the quadratic Rota-Baxter noncommutative pre-Poisson algebra $(A,\succ,\prec,\ast,-\lambda I-P,\omega)$ of non-zero weight $\lambda$.
Proof.
By Proposition 4.10 and Theorem 5.10, there is a quadratic Rota-Baxter noncommutative pre-Poisson algebra $(A,\succ,\prec,\ast,P^{\prime},\omega^{\prime})$ of non-zero weight $\lambda$ corresponds to the factorizable noncommutative pre-Poisson bialgebra $(A,\succ,\prec,\ast,\Delta_{\succ,\tau(r)},\Delta_{\prec,\tau(r)},\delta_{\tau(r)})$. Using Theorem 5.14,
$\omega^{\prime}(x,y)=\lambda\langle T_{\tau(r)-r}^{-1}(x),y\rangle=-\omega(x,y).$
Using Eqs. (143) and (148),
$P^{\prime}(x)=-T_{\tau(r)}{\omega^{\prime}}^{\sharp}(x)=T_{\tau(r)}\omega^{\sharp}(x)$ $=\lambda T_{\tau(r)}T_{r-\tau(r)}^{-1}(x)=\lambda(T_{r}-T_{r-\tau(r)})T_{r-\tau(r)}^{-1}(x)$
$=\lambda T_{r}T_{r-\tau(r)}^{-1}(x)-\lambda(x)=T_{r}\omega^{\sharp}(x)-\lambda(x)=-P(x)-\lambda(x).$
Thus, the factorizable noncommutative pre-Poisson bialgebra $(A,\succ,\prec,\ast,\Delta_{\succ,\tau(r)},\Delta_{\prec,\tau(r)},\delta_{\tau(r)})$ yields a quadratic Rota-Baxter noncommutative pre-Poisson algebra $(A,\succ,\prec,\ast,-\lambda I-P,-\omega)$ of non-zero weight $\lambda$. Analogously, the converse part holds. ∎
Acknowledgments
This work was supported by the Natural Science Foundation of Zhejiang Province of China (LY19A010001), the Science and Technology Planning Project of Zhejiang Province (2022C01118).
Statements and Declarations
All datasets underlying the conclusions of the paper are available to readers. No conflict of interest exits in the submission of this manuscript.

### v2

@@
 Furthermore, we explicitly construct a correspondence linking factorizable noncommutative pre-Poisson bialgebras to quadratic Rota-Baxter noncommutative pre-Poisson algebras.
-1.1. (Noncommutative) Poisson algebras and (noncommutative) pre-Poisson algebras
+$1.$ (Noncommutative) Poisson algebras and (noncommutative) pre-Poisson algebras
 A Poisson algebra is an algebraic structure that combines a Lie algebra with a commutative associative algebra, linked together by the Leibniz rule.
@@
 Noncommutative Poisson algebras have been studied by many authors from different aspects, see [15, 16, 17, 33, 19] and reference therein.
-Definition 1.1.
+Definition $1.$
 A noncommutative Poisson algebra is a triple $(A,\cdot,[\ ,\ ])$ such that $(A,\cdot)$ is an associative algebra (not necessarily commutative), (A,[,])(A,[\ ,\ ]) is a Lie algebra and the following Leibniz rule holds:
@@
 2.1. (Quasi-) representations of noncommutative pre-Poisson algebras
-Let $A$ be a vector space with a bilinear product $\ast:A\otimes A\longrightarrow A$, $(A,\ast)$ is called a pre-Lie algebra if
+Let AA be a vector space with a bilinear product $\ast:A\otimes A\longrightarrow A$, $(A,\ast)$ is called a pre-Lie algebra if
 $(x\ast y)\ast z-x\ast(y\ast z)=(y\ast x)\ast z-y\ast(x\ast z),~\forall~x,y,z\in A.$
 Definition 2.1.
-A dendriform algebra is a vector space $A$ together with two bilinear maps $\succ,\prec:A\otimes A\longrightarrow A$ such that the following equalities hold for all $x,y,z\in A$:
+A dendriform algebra is a vector space AA together with two bilinear maps $\succ,\prec:A\otimes A\longrightarrow A$ such that the following equalities hold for all $x,y,z\in A$:
 $(x\prec y)\prec z=x\prec(y\succ z+y\prec z),$
@@
 $x\succ(y\succ z)=(x\succ y+x\prec y)\succ z.$
-Definition 2.2.
+Definition $2.$
 A noncommutative pre-Poisson algebra is a quadruple $(A,\succ,\prec,\ast)$ such that $(A,\succ,\prec)$ is a dendriform algebra and $(A,\ast)$ is a pre-Lie algebra satisfying the following compatibility conditions:
@@
 Example 2.3.
-Let $A$ be a 2-dimensional vector space with a basis $\{e_{1},e_{2}\}$. Define bilinear maps $\succ,\prec,\ast:A\otimes A\longrightarrow A$ respectively by (only non-zero multiplications are listed):
+Let AA be a 2-dimensional vector space with a basis $\{e_{1},e_{2}\}$. Define bilinear maps $\succ,\prec,\ast:A\otimes A\longrightarrow A$ respectively by (only non-zero multiplications are listed):
 $e_{1}\succ e_{2}=e_{2},\ \ \ e_{1}\succ e_{1}=e_{1},\ \ \ e_{1}\prec e_{2}=-e_{2},$
@@
 Example 2.4.
-Let $A$ be a 3-dimensional vector space with a basis $\{e_{1},e_{2},e_{3}\}$.
+Let AA be a 3-dimensional vector space with a basis $\{e_{1},e_{2},e_{3}\}$.
 Define binary products $\succ,\prec,\ast:A\otimes A\longrightarrow A$ respectively by (only non-zero multiplications are listed):
@@
 (17) $r_{\cdot}(x)(l_{\ast}(y)-r_{\ast}(y))+(l_{\ast}(y)-r_{\ast}(y))r_{\cdot}(x)=r_{\cdot}(y\ast x-x\ast y).$
-Let $A$ and $V$ be vector spaces.
+Let AA and $V$ be vector spaces.
 For a linear map $f:A\longrightarrow\hbox{End}(V)$, define a linear map $f^{*}:A\longrightarrow\hbox{End}(V^{*})$ by $\langle f^{*}(x)u^{*},v\rangle=\langle u^{*},f(x)v\rangle$ for all $x\in A,u^{*}\in V^{*},v\in V$, where $\langle\ ,\ \rangle$ is the usual pairing between $V$ and $V^{*}$.
@@
   3. (c)
-    If $(V,l_{\succ},r_{\succ},l_{\prec},r_{\prec},l_{\ast},r_{\ast})$ is a quasi-representation of $(A,\succ,\prec,\ast)$, then both $(V,l_{\succ},r_{\prec},-l_{\ast})$ and $(V,l_{\cdot},r_{\cdot},r_{\ast}-l_{\ast},r_{\ast})$ are quasi-representations of the sub-adjacent noncommutative Poisson algebra $(A,\cdot,[\ ,\ ])$, where $r_{\cdot}=r_{\succ}+r_{\prec},~l_{\cdot}=l_{\succ}+l_{\prec}$.
+    If $(V,l_{\succ},r_{\succ},l_{\prec},r_{\prec},l_{\ast},r_{\ast})$ is a quasi-representation of $(A,\succ,\prec,\ast)$, then both $(V,l_{\succ},r_{\prec},l_{\ast})$ and $(V,l_{\cdot},r_{\cdot},r_{\ast}-l_{\ast},r_{\ast})$ are quasi-representations of the sub-adjacent noncommutative Poisson algebra $(A,\cdot,[\ ,\ ])$, where $r_{\cdot}=r_{\succ}+r_{\prec},~l_{\cdot}=l_{\succ}+l_{\prec}$.
 Proof.
@@
   4. (d)
-    $(A,L_{\succ},R_{\prec},-L_{\ast})$ is a representation of the sub-adjacent noncommutative Poisson algebra $(A,\cdot,[\ ,\ ])$ and $(A^{*},R_{\prec}^{*},L_{\succ}^{*},-L_{\ast}^{*})$ is a quasi-representations of $(A,\cdot,[\ ,\ ])$.
+    $(A,L_{\succ},R_{\prec},L_{\ast})$ is a representation of the sub-adjacent noncommutative Poisson algebra $(A,\cdot,[\ ,\ ])$ and $(A^{*},R_{\prec}^{*},L_{\succ}^{*},-L_{\ast}^{*})$ is a quasi-representations of $(A,\cdot,[\ ,\ ])$.
   5. (e)
@@
 (18)-(20), then $A_{1}\ltimes A_{2}$ and $A_{1}\ltimes A_{2}^{*}$ both are coherent noncommutative pre-Poisson algebras.
-2.2. Phase spaces of noncommutative Poisson algebras
+$2.$ Phase spaces of noncommutative Poisson algebras
 We now introduce the notions of symplectic noncommutative Poisson algebras and the phase spaces of noncommutative Poisson algebras.
@@
 Let $(A,\succ,\prec,\ast)$ be the 2-dimensional noncommutative pre-Poisson algebra with a basis $\{e_{1},e_{2}\}$ given in Example 2.3.
-Define a nondegenerate skew-symmetric bilinear form $\omega$ on $A$ as follows: $\omega(e_{1},e_{2})=-\omega(e_{2},e_{1})=1$ and $\omega(e_{2},e_{2})=\omega(e_{1},e_{1})=0$.
+Define a nondegenerate skew-symmetric bilinear form $\omega$ on AA as follows: $\omega(e_{1},e_{2})=-\omega(e_{2},e_{1})=1$ and $\omega(e_{2},e_{2})=\omega(e_{1},e_{1})=0$.
 Then $(A,\succ,\prec,\ast,\omega)$ is a quadratic noncommutative pre-Poisson algebra.
@@
 Definition 2.19.
-A Manin triple of noncommutative pre-Poisson algebras is a quadruple $(A,A_{1},A_{2},\omega)$, where $(A,\omega)$ is a quadratic noncommutative pre-Poisson algebra, $(A_{1},\succ_{1},\prec_{1},\ast_{1})$ and $(A_{2},\succ_{2},\prec_{2},\ast_{2})$ are noncommutative pre-Poisson subalgebras of $A$, such that
+A Manin triple of noncommutative pre-Poisson algebras is a quadruple $(A,A_{1},A_{2},\omega)$, where $(A,\omega)$ is a quadratic noncommutative pre-Poisson algebra, $(A_{1},\succ_{1},\prec_{1},\ast_{1})$ and $(A_{2},\succ_{2},\prec_{2},\ast_{2})$ are noncommutative pre-Poisson subalgebras of AA, such that
   1. (a)
@@
 Proposition 2.20.
-Let $(A,A_{1},A_{2},\omega)$ be a Manin triple of noncommutative pre-Poisson algebras. Then the noncommutative pre-Poisson algebras $A$, $A_{1}$ and $A_{2}$ are coherent.
+Let $(A,A_{1},A_{2},\omega)$ be a Manin triple of noncommutative pre-Poisson algebras. Then the noncommutative pre-Poisson algebras AA, $A_{1}$ and $A_{2}$ are coherent.
 Proof.
@@
 $=$ $0,$
-which indicates that Eq. (6) holds. Thus, $A_{1}$ is coherent. Analogously, $A$ and $A_{2}$ are coherent. ∎
+which indicates that Eq. (6) holds. Thus, $A_{1}$ is coherent. Analogously, AA and $A_{2}$ are coherent. ∎
 Theorem 2.21.
@@
   2. (b)
-    $(A\oplus A^{*},A,A^{*},\omega)$ is a Manin triple of noncommutative pre-Poisson algebras with the bilinear form given by Eq. (66) and the isotropic subalgebras are $A$ and $A^{*}$.
+    $(A\oplus A^{*},A,A^{*},\omega)$ is a Manin triple of noncommutative pre-Poisson algebras with the bilinear form given by Eq. (66) and the isotropic subalgebras are AA and $A^{*}$.
   3. (c)
@@
 In view of Corollary 2.18, $(A,\succ|_{A},\prec|_{A},\ast|_{A})$ and $(A^{*},\succ|_{A^{*}},\prec|_{A^{*}},\ast|_{A^{*}})$ are subalgebras of $(A\oplus A^{*},\succ,\prec,\ast)$.
-It is obvious that $A$ and $A^{*}$ are isotropic.
+It is obvious that AA and $A^{*}$ are isotropic.
 Thus, $(A\oplus A^{*},A,A^{*},\omega)$ is a Manin triple of noncommutative pre-Poisson algebras.
@@
   1. (a)
-    A dendriform coalgebra is a triple $(A,\Delta_{\succ},\Delta_{\prec})$, where $A$ is a vector space and $\Delta_{\succ},\Delta_{\prec}:A\longrightarrow A\otimes A$ are linear maps such that the following conditions hold:
+    A dendriform coalgebra is a triple $(A,\Delta_{\succ},\Delta_{\prec})$, where AA is a vector space and $\Delta_{\succ},\Delta_{\prec}:A\longrightarrow A\otimes A$ are linear maps such that the following conditions hold:
     $(\Delta_{\succ}\otimes I)\Delta_{\prec}=(I\otimes\Delta_{\prec})\Delta_{\succ},\ \ \ (I\otimes\Delta_{\succ})\Delta_{\succ}=(\Delta\otimes I)\Delta_{\succ},\ \ \ (\Delta_{\prec}\otimes I)\Delta_{\prec}=(I\otimes\Delta)\Delta_{\prec},$
@@
 where $Q(x)=I\otimes L_{\succ}(x)-R_{\prec}(x)\otimes I$.
-Proposition 3.3.
+Proposition $3.$
 Let $(A,\succ,\prec)$ be a dendriform algebra and $r=\sum\limits_{i}a_{i}\otimes b_{i}\in A\otimes A$ symmetric.
@@
   1. (a)
-    A pre-Lie coalgebra is a vector space $A$ together with a linear map $\delta:A\longrightarrow A\otimes A$ such that
+    A pre-Lie coalgebra is a vector space AA together with a linear map $\delta:A\longrightarrow A\otimes A$ such that
     $(I\otimes\delta)\delta-(\tau\otimes I)(I\otimes\delta)\delta=(\delta\otimes I)\delta-(\tau\otimes I)(\delta\otimes I)\delta.$
@@
 for all $x,y\in A,a,b\in A^{*}$. $(D=A\oplus A^{*},\succ_{D},\prec_{D},\ast_{D})$ is called the double noncommutative pre-Poisson algebra.
-3.3. Coboundary noncommutative pre-Poisson bialgebras and the noncommutative pre-Poisson Yang-Baxter equation
+$3.$ Coboundary noncommutative pre-Poisson bialgebras and the noncommutative pre-Poisson Yang-Baxter equation
 In this section, we consider the coboundary noncommutative pre-Poisson bialgebras.
@@
 (67)-(68) and (78) respectively.
-For a vector space $A$, the isomorphism $A\otimes A^{*}\simeq Hom(A^{*},A)$ identifies an element $r\in A\otimes A$ with a map $T_{r}:A^{*}\longrightarrow A$. Explicitly,
+For a vector space AA, the isomorphism $A\otimes A^{*}\simeq Hom(A^{*},A)$ identifies an element $r\in A\otimes A$ with a map $T_{r}:A^{*}\longrightarrow A$. Explicitly,
 (107) $T_{r}:A^{*}\longrightarrow A,\ \ \ \langle T_{r}(\zeta),\eta\rangle=\langle r,\zeta\otimes\eta\rangle,\ \ \ \forall~\zeta,\eta\in A^{*}.$
@@
 (18)-(20).
-Assume that $(V^{*},r_{\cdot}^{*},-l_{\prec}^{*},-r_{\succ}^{*},l_{\cdot}^{*},r_{\ast}^{*}-l_{\ast}^{*},r_{\ast}^{*})$ is the dual representation of $A$ given by Proposition 2.6.
+Assume that $(V^{*},r_{\cdot}^{*},-l_{\prec}^{*},-r_{\succ}^{*},l_{\cdot}^{*},r_{\ast}^{*}-l_{\ast}^{*},r_{\ast}^{*})$ is the dual representation of AA given by Proposition 2.6.
 Denote $\hat{A}=A\ltimes V^{*}$ and suppose that $T:V\longrightarrow A$ is a linear map which is identifies an element in $\hat{A}\otimes\hat{A}$ through ($Hom(V,A)\simeq A\otimes V^{*}\subseteq\hat{A}\otimes\hat{A}$).
@@
 (122) $L_{\ast}^{*}(T_{r-\tau(r)}(\zeta))\eta=-L_{\ast}^{*}(T_{r-\tau(r)}(\eta))\zeta,\ \ \ R_{\prec}^{*}(T_{r-\tau(r)}(\zeta))\eta=L_{\succ}^{*}(T_{r-\tau(r)}(\eta))\zeta.$
-Theorem 4.4.
+Theorem $4.$
 Let $(A,\succ,\prec,\ast)$ be a coherent noncommutative pre-Poisson algebra and $r=\sum_{i}a_{i}\otimes b_{i}\in A\otimes A$.
@@
 Let $(A,\succ,\prec,\ast,\Delta_{\succ,r},\Delta_{\prec,r},\delta_{r})$ be a noncommutative pre-Poisson bialgebra.
-Assume that $\{e_{1},...,e_{n}\}$ is a basis of $A$ and $\{e^{*}_{1},...,e^{*}_{n}\}$ is the dual basis.
+Assume that $\{e_{1},...,e_{n}\}$ is a basis of AA and $\{e^{*}_{1},...,e^{*}_{n}\}$ is the dual basis.
 Let
@@
 We then establish the relationship between factorizable noncommutative pre-Poisson bialgebras and these quadratic Rota-Baxter noncommutative pre-Poisson algebras.
-Definition 5.5.
+Definition $5.$
 Assume that $(A,\succ,\prec,\ast,P)$ is a Rota-Baxter noncommutative pre-Poisson algebra of weight $\lambda$ and $(A,\succ,\prec,\ast,\omega)$ a quadratic noncommutative pre-Poisson algebra.
@@
 Then $(A,\cdot,[\ ,\ ],P,\omega)$ is a symplectic Rota-Baxter noncommutative Poisson algebra of weight $\lambda$, where $x\cdot y=x\succ y+x\prec y,~[x,y]=x\ast y-y\ast x$.
-Let $\omega$ be a non-degenerate bilinear form on a vector space $A$. Then there is an isomorphism $\omega^{\sharp}:A\longrightarrow A^{*}$ given by
+Let $\omega$ be a non-degenerate bilinear form on a vector space AA. Then there is an isomorphism $\omega^{\sharp}:A\longrightarrow A^{*}$ given by
 (141) $\omega(x,y)=\langle\omega^{\sharp}(x),y\rangle,~~\forall~x,y\in A.$
@@
 Lemma 5.10.
-Let $(A,\succ,\prec,\ast)$ be a noncommutative pre-Poisson algebra and $\omega$ be a non-degenerate bilinear form on $A$.
+Let $(A,\succ,\prec,\ast)$ be a noncommutative pre-Poisson algebra and $\omega$ be a non-degenerate bilinear form on AA.
 Then $(A,\succ,\prec,\ast,\omega)$ is a quadratic noncommutative pre-Poisson algebra if and only if the corresponding $r_{\omega}\in A\otimes A$ given by Eq.
@@
 Lemma 5.12.
-Let $A$ be a vector space and $\omega$ be a non-degenerate skew-symmetric bilinear form. Let $r\in A\otimes A$, $\lambda\in k$ and a linear map $P$ given by Eq. (143). Then $r$ satisfies
+Let AA be a vector space and $\omega$ be a non-degenerate skew-symmetric bilinear form. Let $r\in A\otimes A$, $\lambda\in k$ and a linear map $P$ given by Eq. (143). Then $r$ satisfies
 (147) $r-\tau(r)=\lambda r_{\omega}$